### v2

# Multigranular Evaluation for Brain Visual Decoding

Abstract
Existing evaluation protocols for brain visual decoding predominantly rely on coarse metrics that obscure inter-model differences, lack neuroscientific foundation, and fail to capture fine-grained visual distinctions. To address these limitations, we introduce BASIC, a unified, multigranular evaluation framework that jointly quantifies structural fidelity, inferential alignment, and contextual coherence between decoded and ground-truth images. For the structural level, we introduce a hierarchical suite of segmentation-based metrics, including foreground, semantic, instance, and component masks, anchored in granularity-aware correspondence across mask structures. For the semantic level, we extract structured scene representations encompassing objects, attributes, and relationships using multimodal large language models, enabling detailed, scalable, and context-rich comparisons with ground-truth stimuli. We benchmark a diverse set of visual decoding methods across multiple stimulus-neuroimaging datasets within this unified evaluation framework. Together, these criteria provide a more discriminative, interpretable, and comprehensive foundation for evaluating brain visual decoding methods.
Codehttps://github.com/weihaox/BASIC
Introduction
Recent advances in brain visual decoding (Takagi and Nishimoto 2023a; Ozcelik and VanRullen 2023; Scotti et al. 2023; Xia et al. 2024a) have achieved remarkable success in reconstructing visual stimuli from the neural activations. However, the evaluation protocols commonly used in this field remain limited in several critical aspects. First, current metrics often saturate across state-of-the-art models, limiting their discriminative capacity and obscuring substantive differences in decoded results. Second, these metrics often lack a neuroscientific foundation, failing to capture the perceptual validity of decoded outputs and their alignment with human-like perception. Third, prevailing evaluation strategies typically fail to reflect the multilevel and structured nature of visual perception, neglecting key components such as object semantics, scene understanding, and contextual reasoning. These limitations hinder rigorous benchmarking of brain decoding models and obscure the specific dimensions along which reconstructions succeed or fall short.
What should brain decoding recover.
Brain visual decoding aims to reconstruct visual experiences from neural activations, recovering not only the appearance of stimuli but also their structure, semantics, and perceptual salience. A well-decoded reconstruction should reflect what the subject consciously perceived, preserving salient objects, their attributes, spatial configuration, and overall scene coherence. Since human visual perception is shaped by attention, context, and prior knowledge, brain decoding must align with the hierarchical nature of vision, spanning from low-level pixel patterns to high-level semantic understanding. Therefore, effective brain visual decoding requires both perceptual accuracy and semantic integrity. It should faithfully capture salient elements and spatial context in line with the subject’s attention, while maintaining consistent inter-object relationships and scene-level coherence.
What should we measure in brain visual decoding.
Current evaluation protocols, such as pixel-wise correlation or feature-based similarity, often fall short in capturing the full complexity of brain visual decoding. Low-level metrics tend to overlook scene semantics and perceptual plausibility, while high-level black-box measures conflate multiple alignment aspects into a single score, offering limited diagnostic insight. They struggle to determine whether the “decoded” details truly originates from brain signals, or if they are instead hallucinated constructs based on prototypical co-occurrences from pretrained generative models conditioned on scene or object labels. Moreover, existing metrics are often saturated, assigning uniformly high scores across diverse methods and thus failing to capture fine-grained distinctions in reconstruction quality. We argue that an effective brain visual decoding metric should be neuroscientifically interpretable, grounded in principles of human visual perception, and meet three key desiderata. First, it should be multigranular, capturing perceptual alignment across multiple abstraction levels – from basic object identification and segmentation to semantic and spatial reasoning about attributes and interactions. Second, it should be semantically aligned with human perception, reflecting the way humans interpret scenes, objects, and their relationships in a coherent and meaningful manner. Third, it should be diagnostically informative, offering interpretable feedback on what is correct, what is missing, and where the reconstruction fails. Specifically, it should localize and characterize semantic and structural errors in decoded outputs, such as object misidentification, incorrect attributes, or implausible interactions.
Our method: BASIC.
To bridge these gaps, we introduce BASIC (Brain-Aligned Structural, Inferential, and Contextual similarity), a unified, multigranular evaluation framework for brain visual decoding. BASIC integrates structure matching and semantic reasoning to systematically quantify structural, inferential, and contextual alignment between decoded results and reference stimuli across diverse brain-to-vision tasks in a structured and interpretable manner.
BASIC decomposes evaluation into three complementary perspectives, each reflecting a core aspect of perceptual alignment in brain visual decoding: (a) structural similarity quantifies the reconstructed visual structures, capturing spatial organization and categorical boundaries. This consistency with reference stimuli is operationalized through a granularity-aware mask correspondence across the foreground, semantic, instance, and component levels; (b) inferential similarity measures semantic accuracy, evaluating whether the decoded image conveys the same entities and conceptual content as the reference. This is computed via structured comparisons across object categories, attributes, and inter-object relational graphs extracted using captions from multimodal large language models; (c) contextual similarity assesses perceptual and cognitive plausibility, examining whether the reconstructed scene forms an internally coherent and contextually appropriate whole. This is evaluated using MLLM-based scene reasoning to quantify narrative consistency and global scene coherence.
BASIC offers a comprehensive view of decoding performance across modalities (image, video, 3D) and neuroimaging types (fMRI, EEG). Our framework facilitates both quantitative comparisons and qualitative diagnostics, allowing for fine-grained benchmarking of brain decoding models across datasets. BASIC aims to (1) offer a more detailed and interpretable evaluation of brain decoding results, (2) quantify semantic plausibility in terms more closely aligned with human cognitive processes, and (3) facilitate comparison of decoding methods within a unified evaluation framework applicable across diverse stimulus-neuroimaging datasets. We hope our method contributes to establishing a more systematic foundation for brain visual decoding evaluation.
Figure 1: BASIC evaluates decoded reconstructions along two axes: high-level semantic (BASIC-H) and low-level structural (BASIC-L) similarities. For the semantic axis (inferential and contextual), we extract and compare structured representations from reconstructed and ground-truth images. For the structural axis, we compute mask-based matching across fine-grained segmentation types of identified scenes and objects: salient, semantic, instance, and parts.
Related Work
Brain visual decoding.
The task of brain visual decoding aims to reconstruct perceived visual stimuli, such as images, videos, or 3D shapes, from recorded neural activations. Recent progress in this domain has been closely tied to advancements in computational modeling frameworks. Early methods primarily involved training neural networks from scratch to learn mappings between brain activity and visual features. However, these models often suffered from limited fidelity and exhibited artifacts and poor visual quality in the reconstructed outputs. Recent developments have led to significant improvements due to the emergence of multimodal generative models (Radford et al. 2021; Rombach et al. 2022; Xu et al. 2023) and large-scale brain-stimulus datasets (Wen et al. 2018; Allen et al. 2022; Gao et al. 2024; Guo et al. 2025). These resources have facilitated a new generation of decoding paradigms that leverage intermediate representations from pretrained generative models and align them with neural responses through various strategies, including linear regression (Ozcelik and VanRullen 2023; Takagi and Nishimoto 2023a), diffusion priors (Scotti et al. 2023, 2024), or feature-wise reconstruction (Xia et al. 2024a; Xia and Öztireli 2025b, a). Recent efforts have also focused on eliminating dependence on subject-specific encoders (Scotti et al. 2024; Wang et al. 2024a; Xia et al. 2024a; Tian et al. 2025; Gong et al. 2025), demonstrating promising progress toward subject-independent brain decoding. Besides the commonly used fMRI-image Natural Scenes Dataset (NSD) (Allen et al. 2022), a growing body of research has explored alternative combinations of neuroimaging modalities and stimuli, such as fMRI-video (Wen et al. 2018), fMRI-3D (Gao et al. 2024), EEG-image (Grootswagers et al. 2022), EEG-video (Liu et al. 2024b), and EEG-3D (Guo et al. 2025) decoding. These explorations broaden the scope of brain decoding and offer new insights into the neural representation of dynamic and immersive visual experiences.
Brain decoding evaluation.
The evaluation metrics for visual brain decoding lack a universally accepted standard. Different stimulus-neuroimaging combinations employ varying evaluation protocols. For instance, the following eight metrics are commonly used on NSD: PixCorr, SSIM (Wang et al. 2004), AlexNet (Krizhevsky, Sutskever, and Hinton 2017)-2/5, Inception (Szegedy et al. 2016), CLIP (Radford et al. 2021), EffNet (Tan and Le 2019), and SwAV (Caron et al. 2020). The first four metrics are considered low-level, focusing on perceptual similarity and pixel-wise or structural correspondence. In contrast, the latter four are used to evaluate semantic decoding or high-level representations, emphasizing the overall theme or content of the reconstructed images. For other stimulus-neuroimaging setups, evaluation protocols remain inconsistent. Metrics such as $n$-way classification accuracy (Guo et al. 2025), CLIP-based Pearson correlation (Gong et al. 2024), mask-matching ratios (Li et al. 2025), and modality-specific evaluation metrics (Gao et al. 2025) have all been reported in the recent literature. However, prior metrics overlook the hierarchical nature of perception and struggle to distinguish models with subtle differences. In contrast, our BASIC framework offers a unified, multigranular evaluation across structural, inferential, and contextual dimensions, enabling more nuanced and interpretable comparisons.
BASIC: Evaluating Brain Visual Decoding
BASIC provides a versatile evaluation framework for brain visual decoding across diverse combinations of visual stimuli and neuroimaging modalities. This section details the evaluation dimensions and the two complementary modules of our BASIC metric: BASIC-H integrates the inferential and contextual dimensions of BASIC into a unified measure of high-level semantic correspondence, while BASIC-L quantifies low-level structural alignment.
Evaluation Dimension
To gain deeper insight into how brain decoding models represent visual semantics across different levels of abstraction, we develop a structured evaluation protocol grounded in key conceptual dimensions of human visual perception: scenes, objects, attributes, and relations. Each dimension is further divided into subcategories as outlined in Table 1.
Scene reflects global properties, including overall layout, geometric structure, event context, and stylistic tone. This probes the holistic configuration and contextual coherence.
Object evaluates recognition and differentiation of objects in the scene, including both categorical accuracy (e.g., “cat” vs. “dog”) and semantic granularity—from object universality (e.g., identifying any “car”) to specificity (e.g., distinguishing an “ambulance” from a “sedan”).
Attribute covers appearance cues (e.g., color, texture, material) and spatial properties (e.g., location, count). We also extract symbolic visual information such as text using optical character recognition.
Relation assesses interactions between entities, including inter-object and object-scene relationships, as well as spatial and part-whole relations (e.g., “wheel is part of car”), physical interactions (e.g., “man holding umbrella”), and dynamic cues such as posture, expression, and motion.
Camera captures the photographer’s perspective and natural conditions during the shot, including viewpoint (e.g., frontal, top-down), motion (e.g., zoom, pan), and lighting.
These dimensions are motivated by principles in visual neuroscience and cognitive psychology (Desimone et al. 1984; Puce et al. 1996), and aligned with the structure of scene understanding in large multimodal models (Dong et al. 2024; Lu et al. 2025; Liu et al. 2023). These dimensions are then organically integrated into structural, inferential, and contextual similarity components within our framework, converted into two sub-indicators, as outlined below.
Table 1: Evaluation dimension and details.
Dims. Details
scene layout, geometry, event, action, style
object category, commonality (universality), specificity (differentiation)
attribute appearance (color, texture, material), position, quantity, symbols and text
relation spatial or part-whole relation, interaction, kinematics (motion, speed)
camera lighting, camera angle, camera motion
Table 2: BASIC-H scores across stimulus-neuroimaging datasets. This metric, with sub-indicators for precision, recall, and F1 – evaluated on objects, attributes, and relations – quantifies high-level semantic correspondence between reconstructions and stimuli. Best and second best are highlighted.
Method Object Attribute Relation BASIC-H
P R F1 P R F1 P R F1
NSD (Allen et al. 2022)
SDRecon (Takagi and Nishimoto 2023a) 55.59 53.05 53.79 10.12 38.73 14.96 40.15 38.71 39.06 35.31
BrainDiffuser (Ozcelik and VanRullen 2023) 57.87 59.11 58.09 13.35 45.82 19.43 43.20 44.42 43.50 39.71
MindEye (Scotti et al. 2023) 62.94 60.64 61.26 18.14 51.17 25.06 49.98 48.42 48.84 44.30
DREAM (Xia et al. 2024b) 65.63 63.06 63.56 18.97 50.68 25.92 53.45 53.21 52.91 46.37
MindEye2 (Scotti et al. 2024) 62.57 62.12 61.72 17.86 50.16 24.71 49.72 49.17 49.07 44.39
MindBridge (Wang et al. 2024a) 59.00 58.35 58.19 13.70 45.69 19.49 45.75 45.78 45.43 40.16
UMBRAE (Xia et al. 2024a) 62.00 61.86 61.44 17.51 51.32 24.49 48.91 48.71 48.45 44.06
NeuroPictor (Huo et al. 2024) 63.00 61.05 61.38 17.92 50.13 24.66 49.62 49.08 48.98 44.21
NeuroVLA (Shen et al. 2024) 65.36 65.03 64.57 21.27 53.73 28.65 53.80 52.86 52.95 47.88
SepBrain (Wang et al. 2024b) 62.10 60.19 60.57 16.62 48.71 23.31 48.03 47.55 47.44 43.04
UniBrain (Wang et al. 2024b) 59.03 58.21 58.07 13.27 43.89 19.02 45.55 45.58 45.25 39.89
STTM (Liu et al. 2025) 64.50 62.50 62.88 19.53 51.70 26.64 51.17 50.28 50.36 45.88
MindTuner (Gong et al. 2025) 62.55 62.64 61.95 18.06 49.18 24.73 50.22 50.10 49.80 44.63
BrainGuard (Tian et al. 2025) 63.63 62.36 62.43 18.66 52.40 25.84 51.25 50.72 50.60 45.43
EEG-Things  (Grootswagers et al. 2022)
ATM (Li et al. 2024) 44.54 44.51 43.77 9.65 38.48 14.46 36.74 36.63 36.28 30.55
CognitionCapturer (Zhang et al. 2025) 42.20 43.99 42.59 7.56 34.56 11.67 34.31 35.74 34.70 28.64
CC2017 (Wen et al. 2018)
MinD-Video (Chen, Qing, and Zhou 2023) 47.56 45.05 45.34 6.95 29.27 10.57 32.42 31.10 31.32 28.63
NeuroClips (Gong et al. 2024) 62.82 61.68 61.28 17.12 50.82 24.30 53.91 56.33 54.42 45.12
DecoFuse (Li et al. 2025) 47.35 47.12 46.74 6.26 28.12 9.77 32.75 32.18 32.11 29.03
SEED-DV (Liu et al. 2024b)
EEG2Video (Liu et al. 2024b) 66.05 65.75 64.36 25.43 51.85 32.20 54.63 56.02 54.59 49.54
fMRI-Shape (Gao et al. 2024)
MinD-3D (Gao et al. 2024) 41.52 36.42 37.59 13.52 44.00 19.15 42.49 41.12 41.26 30.95
MinD-3D++ (Gao et al. 2025) 41.35 37.92 38.12 20.03 61.89 27.80 44.82 43.39 43.57 35.08
EEG-3D (Guo et al. 2025)
Neuro-3D (Guo et al. 2025) 35.79 32.86 34.26 6.32 27.92 10.32 23.41 29.87 26.24 23.08
Table 3: BASIC-L scores across datasets. This metric evaluates low-level structural correspondence between reconstructed and reference images at four granularities: foreground saliency, semantic consistency, instance separation, and part-level delineation.
Method F (Foreground) B (Binary) S (Semantic) I (Instance) P (Part) BASIC-L
IoU AP IoU AP IoU AP IoU AP IoU AP
NSD (Allen et al. 2022)
SDRecon (Takagi and Nishimoto 2023b) 9.03 13.06 38.90 49.97 15.08 21.43 17.84 1.07 4.38 6.79 11.81
BrainDiffuser (Ozcelik and VanRullen 2023) 17.96 20.21 38.98 45.85 18.66 20.78 20.09 1.94 7.86 7.82 16.65
MindEye (Scotti et al. 2023) 19.20 22.79 45.53 55.17 18.73 21.94 20.36 1.99 7.49 8.26 17.03
DREAM (Xia et al. 2024b) 23.62 26.10 46.03 57.10 21.15 24.13 21.41 2.32 9.22 8.79 19.57
MindEye2 (Scotti et al. 2024) 25.29 26.27 47.93 57.52 24.33 25.68 24.09 3.45 12.32 11.03 22.16
MindBridge (Wang et al. 2024a) 16.24 19.04 40.51 48.95 16.87 19.81 18.61 1.54 6.30 6.78 15.00
UMBRAE (Xia et al. 2024a) 21.33 24.62 40.96 48.53 18.94 21.16 20.29 2.15 8.43 8.47 17.89
NeuroPictor (Huo et al. 2024) 29.45 31.29 47.79 56.48 27.97 29.57 26.47 4.08 17.17 15.84 25.88
NeuroVLA (Shen et al. 2024) 16.03 20.85 44.09 54.42 13.84 17.54 17.57 1.34 4.38 5.57 13.54
SepBrain (Wang et al. 2024b) 21.22 23.72 45.06 53.47 20.88 23.32 21.98 2.51 8.79 8.98 18.84
UniBrain (Wang et al. 2024b) 15.24 18.02 37.48 45.09 14.99 17.75 17.45 1.06 5.54 6.07 13.79
STTM (Liu et al. 2025) 27.31 29.44 49.35 59.05 24.61 26.37 24.19 3.50 12.55 11.65 22.90
MindTuner (Gong et al. 2025) 15.38 16.66 40.74 49.02 21.46 24.50 21.49 1.97 8.16 8.13 16.98
BrainGuard (Tian et al. 2025) 25.90 28.07 49.22 58.99 23.34 25.21 23.98 3.29 10.82 10.32 21.76
EEG-Things  (Grootswagers et al. 2022)
ATM (Li et al. 2024) 13.87 18.77 39.46 50.14 22.85 34.69 22.02 2.15 11.13 17.07 17.60
CognitionCapturer (Zhang et al. 2025) 10.26 12.71 33.31 40.67 21.37 29.84 20.75 1.37 13.08 17.26 16.22
CC2017 (Wen et al. 2018)
MinD-Video (Chen, Qing, and Zhou 2023) 12.82 24.20 44.89 53.50 20.29 33.57 19.87 3.53 6.83 14.69 15.25
NeuroClips (Gong et al. 2024) 24.37 31.59 65.51 73.39 28.23 36.00 28.74 7.73 9.82 13.14 23.52
DecoFuse (Li et al. 2025) 12.32 18.07 49.87 59.46 22.33 34.63 20.87 2.74 4.06 9.80 15.31
SEED-DV (Liu et al. 2024b)
EEG2Video (Liu et al. 2024b) 27.77 32.82 57.26 69.97 22.93 31.96 23.41 3.51 3.10 7.10 20.54
fMRI-Shape (Gao et al. 2024)
MinD-3D (Gao et al. 2024) 5.69 8.10 5.69 8.10 19.91 29.35 19.38 1.45 15.96 25.30 14.72
MinD-3D++ (Gao et al. 2025) 3.80 5.23 3.80 5.23 16.03 35.83 17.96 1.83 14.94 36.19 12.62
EEG-3D (Guo et al. 2025)
Neuro-3D (Guo et al. 2025) 2.39 3.65 2.39 3.65 13.77 25.92 12.25 1.02 12.33 18.56 9.69
BASIC-H
BASIC-H quantifies high-level semantic correspondence by integrating inferential and contextual similarities, which respectively capture complementary aspects of conceptual accuracy and narrative coherence, for a holistic evaluation of reconstructed scene semantics. Specifically, BASIC-H operates in an automated pipeline with three steps: (1) detailed description generation, where state-of-the-art MLLMs produce semantic-rich captions for both reconstructed and reference images; (2) semantic graph construction, which parses captions into semantic graphs representing objects, attributes, and relations; and (3) structured semantic matching, which computes semantic correspondace using symbolic concept matching and embedding-based similarities. See Fig. 1 for the method overview.
Detailed description generation.
To support multigranular evaluation, we first prompt MLLMs to produce semantically rich descriptions from images. This description prompt extracts detailed information across key semantic components corresponding to the semantic dimensions outlined in Table 1, including scene context, objects, object attributes, inter-object relations, and camera viewpoint (when applicable). Recent studies indicate that state-of-the-art MLLMs can generate captions on par with human experts, albeit with occasional hallucinations (Lu et al. 2025; Dong et al. 2024).
Semantic graph construction.
Given detailed captions, we first segment them into individual sentences using NLTK (Loper and Bird 2004). We then extract key visual elements – objects, attributes, and relations – using a T5-based factual parser (Li et al. 2023). To handle objects mentioned across multiple sentences in the caption, we consolidate references to the same object into subcaptions and bind attributes to the correct instances, avoiding indiscriminate merging. For directional relations (e.g., subject-object pairs), we concatenate relevant sentences to maintain relational coherence (Lu et al. 2025; Dong et al. 2024).
Structured semantic matching.
The matching process aligns the extracted objects, attributes, and relations in three steps: (a) exact matching, which identifies direct lexical matches between visual elements from the reconstruction and the reference stimuli; (b) synonym matching, which aligns semantically equivalent terms, such as “building” and “edifice” for objects; “bright blue” and “azure” for attributes; “next to” and “beside” for relations; and (c) semantic matching, which computes cosine similarity for any remaining unmatched elements based on contextual meaning. This multi-step process ensures that all visual elements, whether directly mentioned or linguistically varied, are appropriately matched across decoded and reference images.
Finally, we compute semantic alignment scores based on the matching results. To mitigate the effects of omissions and hallucinations in MLLM outputs, we evaluate precision, recall, and F1 scores for each type of visual element (object, attribute, relation), based on the matching outcomes. The overall BASIC-H score is computed as a weighted sum of the three F1 scores, with weights of $\alpha_{1}$, $\alpha_{2}$, $\alpha_{3}$ assigned to objects, attributes, and relations, respectively. Precision penalizes hallucinations by measuring the proportion of correctly matched elements among all generated ones, whereas recall captures omissions by evaluating how completely the reference content is recovered. Their combination in F1 score enables a balanced assessment of both over- and under-generation issues common in MLLM outputs.
BASIC-L
BASIC-L evaluates the structural correspondence between the reconstructed and reference images across four granularities: foreground saliency, semantic consistency, instance separation, and part-level delineation. The process first extracts key hierarchical visual components and organizes them into natural language expressions in a progressively granular, whole-to-part format using structured prompts. It then applies referring expression comprehension to localize specific objects in the scene based on these expressions. The evaluation follows two steps:
Semantic instance categorization.
We prompt MLLMs to generate structured semantic annotations from images. These annotations include object-centric identifications that specify spatial roles (e.g., foreground or background), semantic categories (e.g., “dog”, “tree”), and part-level component decompositions (e.g., “head”, “leg”). This process incorporates a multi-level decomposition of visual elements, capturing a hierarchical view of the scene components and facilitating further grounded segmentation and multigranular structural alignment.
Progressive granular segmentation.
The categories from these structured annotations are then used as prompts for referring expression comprehension methods (Ren et al. 2023; Kirillov et al. 2023), which produces segmentation masks for each identified object and component. The segmentation results are decided by the text and box thresholds, where only the highest-similarity boxes exceeding the box threshold and words with similarity scores above the text threshold are considered as predicted labels (Liu et al. 2024a).
Hierarchical score calculation.
We compare predicted masks against those from the reference image across four granularities: salient (foreground categories, F), binary (foreground and background categories, B), semantic (all distinct categories, S), instance (individual instances of all identified object categories, I), and part (subobject components, mostly salient foreground objects, P), computing intersection-over-union (IoU) and average precision (AP).
We aggregate saliency, semantic, instance, and part-level segmentation scores into the overall BASIC-L metric via a weighted sum of respective IoUs using weights of $\beta_{1}$, $\beta_{2}$, $\beta_{3}$, $\beta_{4}$ for each granularity. This scheme prioritizes global layout and object-level coherence while accounting for fine-grained details, reflecting the limited granularity of information captured during the brain data acquisition experiments.
Experiments
Experimental Setup
Figure 2: BASIC performance.
Implementation details.
We use LLaVA-1.6-13B (Liu et al. 2023) as the default MLLM for detailed description generation and semantic instance categorization. Ground truth captions are obtained from human experts through error correction, missing element addition, and hallucination removal, based on GPT-4o-generated captions, for better alignment with human judgment. For multigranular segmentation, we employ Grounded-SAM2 (Ren et al. 2023), an open-source referring expression comprehension tool. The default weights for BASIC-H and BASIC-L are set to 4:4:2 and 3:2.5:2.5:2, respectively, reflecting the empirically informed importance of each sub-indicator in capturing semantic and structural alignment. Experiments are conducted on an NVIDIA A100 GPU. To ensure broad applicability and methodological consistency, we refrain from using MLLMs or segmentation methods specifically designed for video or 3D data. Instead, we retain an image-based framework with automatic selection of representative video frames and rendered 3D views that best capture key semantics and structure from stimuli and decoded reconstructions.
Experimental Results
While previous metrics suffer from score saturation and fail to effectively distinguish between method performances (see supplementary material), our metrics enable more nuanced evaluation. They offer clearer differentiation across objects, attributes, and relations for high-level semantics, as well as across salient, semantic, instance, and part-level granularity for low-level structure. Detailed descriptions of each dimension are provided in Table 2 and Table 3.
Semantics: object, attribute, and relation.
Table 2 presents a comparison of object, attribute, and relation prediction performance across several datasets. While prior metrics often produce saturated or undifferentiated scores across models, our proposed multigranular evaluation reveals meaningful variations in object coherence, attribute accuracy, and relational plausibility, offering a more fine-grained diagnostic lens. Here, for methods on NSD (Allen et al. 2022), NeuroVLA (Shen et al. 2024), DREAM (Xia et al. 2024b), and STTM (Liu et al. 2025) achieve the highest BASIC-H scores, reflecting their strength in modeling rich visual semantics. This likely stems from detailed caption generation, complex visual semantics modeling, cross-subject training, or their combinations. A closer comparison between NeuroVLA and DREAM shows that while both models reconstruct correct instances (achieving higher precision) across the three semantic dimensions, NeuroVLA misses fewer relevant instances (better recall). In contrast, SDRecon (Takagi and Nishimoto 2023b), BrainDiffuser (Ozcelik and VanRullen 2023), and UniBrain (Wang et al. 2024b) lag behind with BASIC-H scores. These methods struggle notably with the identification of object categories, attributes, and relations, indicating limitations in handling semantics or generalizing across visual categories.
Structure: salient, semantic, instance, and part.
Table 3 reports performance across salient, binary, semantic, instance, and part segmentation matching scores. For decoding methods on NSD, NeuroPictor (Huo et al. 2024) leads with the highest BASIC-L score, driven by its superior performance in instance and part segmentation. This reflects its ability to reconstruct fine-grained visual details and delineate object boundaries from neural signals. STTM (Liu et al. 2025) and MindEye2 (Scotti et al. 2024) perform competitively. SDRecon (Takagi and Nishimoto 2023b) and NeuroVLA (Shen et al. 2024) show relatively poor spatial structural fidelity. BrainGuard (Tian et al. 2025) achieves competitive scores in salient and semantic categories but performs less satisfactorily in instance and part segmentation. This indicates that while the model captures the overall structure of salient objects and categories, it faces challenges in distinguishing between multiple instances within the same category and identifying sub-component relationships.
Fig. 2 provides performance of visual decoding models on NSD in terms of structural alignment (BASIC-L) and semantic alignment (BASIC-H). Each point represents a model, revealing trade-offs between spatial fidelity and semantic coherence. Models closer to the top-right demonstrate stronger performance across both dimensions. For structural alignment, NeuroPictor (Huo et al. 2024), STTM (Liu et al. 2025), and MindEye2 (Scotti et al. 2024) achieve the highest BASIC-L scores, reflecting better spatial reconstruction. In contrast, semantic alignment (BASIC-H) is best captured by NeuroVLA (Shen et al. 2024), DREAM (Xia et al. 2024b), and STTM (Liu et al. 2025), indicating superior semantic preservation.
Figure 3: BASIC demonstrates stable and consistent performance in method evaluation across variations in (a) MLLMs (Liu et al. 2023), (b) prompting strategies, and (c) thresholds for box and text (Ren et al. 2023).
Figure 4: Qualitative examples with BASIC-H and BASIC-L scores, including sub-indicators.
Discussion
Toward open, stable, and versatile evaluation.
Open: Our evaluation pipeline is designed to be model-agnostic, avoiding reliance on proprietary or task-specific components to ensure broad applicability and reproducibility. While models like GPT-4o are shown to have stronger multimodal performance in captioning benchmarks (Lu et al. 2025; Dong et al. 2024), they are not suitable for open, large-scale benchmarking due to API restrictions and cost. We use LLaVA-1.6-13B, which balances captioning accuracy and computational efficiency. Stable: The methods under our metrics demonstrate stable and consistent performance despite variations in (a) MLLMs, (b) prompting strategies, and (c) thresholds for box and text, as shown in Fig. 3. The relative ranking and discriminative power across decoding methods remain consistent, although absolute scores may vary slightly. The stable BASIC-H results across different MLLMs and prompts indicate that our metric reliably captures semantic performance differences among methods. Similarly, the BASIC-L scores provide consistent structural evaluation across varying text and box threshold settings, where only the highest similarity boxes exceeding the box threshold and words with similarity scores above the text threshold are considered predicted labels. Versatile: For video and 3D data evaluation, instead of using modality-specific tools, we maintain the same image-based pipeline, with automatic selection of representative video frames and 3D-rendered views to ensure compatibility across formats without compromising evaluation quality.
Toward informative diagnostic insight.
This multigranular, interpretable feedback provides fine-grained diagnostic insights into brain-based visual decoding and can help uncover blind spots. Decoding methods that rely on pretrained models benefit from powerful generative capabilities but are also susceptible to systematic biases, often introducing hallucinated details such as prototypical co-occurrence patterns. For instance, savannah may be infered when the decoded scene label is giraffe even though the original stimulus was a zoo. These hallucinations tend to affect high-level contextual information—such as scene types or typical object-environment associations—whereas low-level attributes (e.g., blue) and spatial relationships (e.g., to the left) are less likely to be inferred without support from brain-derived signals. BASIC-H mitigates this confound and allows us to better isolate the contribution of genuine brain-derived information and more accurately assess the true decoding performance. BASIC-H penalizes misidentifications, such as hallucinated objects, incorrect attributes, and implausible relationships. These semantic-level discrepancies are directly reflected in the precision and recall of object categories, attributes, and relations, offering interpretable signals beyond opaque scores from pretrained networks. The segmentation-based BASIC-L scores also serve diagnostic purposes, measuring structure correspondance across four granularities: salient, semantic, instance and part. Fig. 4 presents qualitative examples illustrating how BASIC-H and BASIC-L enable fine-grained breakdowns of semantic and structural alignment. In the top reconstruction example from NeuroVLA (Shen et al. 2024), the concepts of “dog” and “man” are correctly predicted but placed in an incorrect scene. The bottom example from MindEye2 in the airplane scenario correctly predicts both semantic and structural aspects, therefore achieving a high score at both dimensions.
Towards cross-dataset performance dissection.
Our evaluation follows a unified protocol across visual modalities, enabling cross-dataset dissection. The fMRI-based image reconstructions (Ozcelik and VanRullen 2023; Scotti et al. 2023) consistently outperform EEG-based counterparts (Li et al. 2024; Zhang et al. 2025), attributable to intrinsic limitations of EEG in spatial resolution and information. The fMRI-to-video reconstructions (Gong et al. 2024; Chen, Qing, and Zhou 2023; Li et al. 2025) currently struggles to preserve the structure of salient objects, but performs comparably to image modalities in terms of instance semantic identification and structural preservation. The high performance in EEG-to-video decoding (Liu et al. 2024b) appears to reflect the simplicity of stimuli – short clips with prominent foregrounds and minimal background clutter. Future work explore more semantically rich and visually complex video scenarios. For 3D reconstruction, both fMRI-based and EEG-based visual decoding methods (Gao et al. 2024, 2025; Guo et al. 2025) struggle to recover even basic semantics, despite simple, canonical structures of the target object categories. The incapability in semantic identification and structure preservation highlights the need for decoding models with stronger semantic and geometric priors.
Conclusion
We present a brain-based visual decoding evaluation framework that captures the multigranular nature of human visual perception. By integrating mask-based segmentation alignment with structured object-attribute-relation similarity, our approach enhances performance discriminability, neuroscientific validity, and semantic interpretability. The resulting BASIC metrics provide a comprehensive assessment across structural precision, inferential accuracy, and contextual coherence. We benchmark diverse brain visual decoding models across six major stimulus-neuroimaging datasets under a unified evaluation protocol, establishing the first standardized, interpretable, and extensible benchmark for this task.
Acknowledgements. This work was supported by a UKRI Future Leaders Fellowship [grant number G104084].
Appendix
Supplemental Material for “Multigranular Evaluation for Brain Visual Decoding”
Appendix A Details on Stimulus-Neuroimaging Datasets
This section provides details on stimulus-neuroimaging datasets. Table S1 summarizes the core properties of these stimulus-neuroimaging combinations. The Natural Scenes Dataset (Allen et al. 2022) has recently garnered significant attention, with many methods proposed for its analysis, and most of our experiments focus on methods applied to this dataset for comprehensive and detailed analysis.
NSD.
NSD (Allen et al. 2022) is the largest public 7 Tesla fMRI scans, featuring brain recordings from eight participants who passively viewed images from the Common Objects in Context (COCO) dataset (Lin et al. 2014) for up to 40 hours in an MRI machine. Each image was shown for three seconds and repeated three times across 30-40 scanning sessions, resulting in 22,000-30,000 fMRI response trials per participant.
EEG-Things.
EEG-Things (Grootswagers et al. 2022) provides neuroimaging recordings for a systematic collection of objects and concepts, featuring electroencephalography (EEG) responses from 50 subjects to a stimulus set of 1,854 object concepts and 22,248 images.
CC2017.
CC2017 (Wen et al. 2018) is a public benchmark consisting of video clips and corresponding fMRI recordings collected from three subjects using a 3T MRI scanner. The training set includes 18 segments of 8-minute video clips (totaling 2.4 hours), yielding 4,320 paired fMRI-video examples. The test set includes 5 segments (totaling 40 minutes), providing 1,200 samples.
SEED-DV.
SEED-DV (Liu et al. 2024b) includes EEG recordings from 20 subjects using a 62-channel system as they viewed 1,400 video clips covering 40 object concepts. These concepts were grouped into 9 broader categories, offering dynamic visual stimuli paired with corresponding neural responses.
fMRI-3D.
The fMRI-3D dataset comprises two successive components for neural decoding of 3D visual perception. fMRI-Shape (Gao et al. 2024) collects fMRI responses from 14 participants viewing 360-degree videos of 1,624 3D objects spanning 55 categories, totaling 123,200 frames. The subsequent fMRI-Objaverse (Gao et al. 2025) further expands category diversity and object coverage, with data from 5 participants exposed to 3,142 3D objects spanning 117 categories, each accompanied by descriptive text captions.
EEG-3D.
EEG-3D (Guo et al. 2025) provides extensive EEG recordings from 12 participants who viewed 3D objects across 72 categories, rendered as both videos and images. Each participant underwent approximately 5.5 hours of recording. The dataset includes 10 objects per category, selected for shape diversity, with each object paired with a descriptive text caption.
Table S1: Details on stimulus-neuroimaging datasets.
Dataset Venue Neuroimaging Stimulus Method
NSD (Allen et al. 2022) Nature Neuroscience 2022 fMRI Image Takagi and Nishimoto (2023b); Scotti et al. (2023, 2024)
EEG-Things (Grootswagers et al. 2022) Scientific Data 2022 EEG Image Li et al. (2024); Zhang et al. (2025)
CC2017 (Wen et al. 2018) Cerebral Cortex 2018 fMRI Video Li et al. (2025); Gong et al. (2024)
SEED-DV (Liu et al. 2024b) NeurIPS 2024 EEG Video Liu et al. (2024b)
fMRI-3D (Gao et al. 2024) ECCV 2024 fMRI 3D Shape Gao et al. (2024, 2025)
EEG-3D (Guo et al. 2025) CVPR 2025 EEG 3D Shape Guo et al. (2025)
The latter three datasets, SEED-DV (Liu et al. 2024b), fMRI-3D (Gao et al. 2024), EEG-3D (Guo et al. 2025), are the most recently proposed datasets, with limited comparisons available. We share their results to demonstrate that the proposed evaluation framework can be extended to different stimulus-neuroimaging datasets.
Appendix B Details on Prior Evaluation Procedure
Brain visual decoding lacks consistent evaluation standards, with different datasets often employing distinct metrics for performance assessment. This section introduces the common evaluation metrics used in the field and outlines how these metrics assess model performance.
Table S2 presents results on NSD and EEG-Things using eight commonly used metrics: PixCorr, SSIM (Wang et al. 2004), AlexNet (Krizhevsky, Sutskever, and Hinton 2017)-2/5, Inception (Szegedy et al. 2016), CLIP (Radford et al. 2021), EffNet (Tan and Le 2019), and SwAV (Caron et al. 2020). PixCorr measures the Pearson correlation between the pixel values of the ground-truth (GT) image and the reconstruction, quantifying the pixel-level similarity. SSIM (Structural Similarity Index) (Wang et al. 2004) compares the structural and textural features of the GT and the reconstruction. The AlexNet-2/5 metrics refer to the similarity between the embeddings of the GT and the reconstruction, based on the deep features extracted from the second and fifth layers of the pretrained AlexNet (Krizhevsky, Sutskever, and Hinton 2017), respectively. Similarly, Inception refers to the comparison of the final pooling layer in InceptionV3 (Szegedy et al. 2016), and CLIP corresponds to the comparison of embeddings from the final layer of CLIP-Vision (Radford et al. 2021). EffNet-B and SwAV metrics refer to the distance or similarity between the GT embedding and the reconstructed embedding using EfficientNet-B1 (Tan and Le 2019) and SwAV-ResNet50 (Caron et al. 2020), respectively.
Table S3 presents a quantitative evaluation of video reconstruction on CC2017 (Wen et al. 2018), using 2-way and 50-way semantic accuracy as the semantic metrics, and the ratio of foreground-background matching (mask matching ratio, MMR) between the ground truth and decoded images as the spatial-level metric. Table S5, Table S5, and Table S6 shows quantitative evaluation of visual reconstruction on SEED-DV (Liu et al. 2024b), EEG-3D (Guo et al. 2025), and fMRI-3D (Gao et al. 2024), with results excerpted from the corresponding papers. For ease of comparison, we also include BASIC-L and BASIC-H evaluation results. As shown, while NSD metrics saturates in scores and fails to distinguish method performance, our metrics, BASIC-L and BASIC-H, provide clearer differentiation and more discriminative information across objects, attributes, and relations for high-level semantics, as well as salient, semantic, instance, and part granularity for low-level structure. Details for each dimension are in Table 2 and Table 3.
Despite slight variations in specific definitions, these evaluations utilize n-way identification metrics, which measure the percentage of cases where the ground truth embedding or category is closer to its corresponding reconstruction than to other reconstructions. This n-way identification metric is problematic for providing consistent and nuanced assessments in cross-model comparisons, as the simplistic criterion only requires the reconstruction to be closer to the ground truth than a randomly selected alternative and each model is evaluated against different sets of reconstructions.
Other modality-specific metrics are also introduced. For instance, fMRI-3D evaluation uses common 3D reconstruction metrics, including Fréchet Point Cloud Distance (FPD) (scaled by $\times 10^{-1}$), Chamfer Distance (CD) (scaled by $\times 10^{2}$), and Earth Mover’s Distance (EMD) (scaled by $\times 10^{2}$). These metrics are computed by sampling point clouds from both the ground truth and the generated meshes.
Table S2: Quantitative evaluation for fMRI-to-image reconstruction on NSD (Allen et al. 2022) and EEG-to-image reconstruction on EEG-Things (Grootswagers et al. 2022) following standard NSD and our BASIC metrics. While NSD metrics saturates in scores and fails to distinguish method performance, our metrics provide clearer differentiation and more discriminative information. See detailed sub-scores in Table 2 and Table 3.
Method Low-Level High-Level
PixCorr $\uparrow$ SSIM $\uparrow$ Alex(2) $\uparrow$ Alex(5) $\uparrow$ BASIC-L $\uparrow$ Incep $\uparrow$ CLIP $\uparrow$ EffNet $\downarrow$ SwAV $\downarrow$ BASIC-H $\uparrow$
NSD  (Allen et al. 2022)
SDRecon (Takagi and Nishimoto 2023b) 0.246 0.410 78.9% 85.6% 11.81 83.8% 82.1% 0.811 0.504 35.31
BrainDiffuser (Ozcelik and VanRullen 2023) 0.273 0.365 94.4% 96.6% 16.65 91.3% 90.9% 0.728 0.422 39.71
MindEye (Scotti et al. 2023) 0.319 0.360 92.8% 96.9% 17.03 94.6% 93.3% 0.648 0.377 44.30
DREAM (Xia et al. 2024b) 0.274 0.328 93.9% 96.7% 19.57 93.4% 94.1% 0.645 0.418 46.37
MindEye2 (Scotti et al. 2024) 0.322 0.431 96.1% 98.6% 22.16 95.4% 93.0% 0.619 0.344 44.39
MindBridge (Wang et al. 2024a) 0.151 0.263 87.7% 95.5% 15.00 92.4% 94.7% 0.712 0.418 40.16
UMBRAE (Xia et al. 2024a) 0.283 0.341 95.5% 97.0% 17.89 91.7% 93.5% 0.700 0.393 44.06
NeuroPictor (Huo et al. 2024) 0.229 0.375 96.5% 98.4% 25.88 94.5% 93.3% 0.639 0.350 44.21
NeuroVLA (Shen et al. 2024) 0.265 0.357 93.1% 97.1% 13.54 96.8% 97.5% 0.633 0.321 47.88
SepBrain (Wang et al. 2024b) 0.309 0.317 94.2% 97.4% 18.84 94.5% 95.3% 0.656 0.374 43.04
UniBrain (Wang et al. 2024b) 0.155 0.259 87.8% 95.9% 13.79 92.4% 94.0% 0.691 0.407 39.89
STTM (Liu et al. 2025) 0.333 0.334 95.7% 98.5% 22.90 95.8% 95.7% 0.611 0.338 45.88
MindTuner (Gong et al. 2025) 0.322 0.421 95.8% 98.8% 16.98 95.6% 93.8% 0.612 0.340 44.63
BrainGuard (Tian et al. 2025) 0.313 0.330 94.7% 97.8% 21.76 96.1% 96.4% 0.624 0.353 45.43
EEG-Things  (Grootswagers et al. 2022)
ATM (Li et al. 2024) 0.160 0.345 0.776 0.866 17.60 0.734 0.786 - 0.582 30.55
CogCapturer (Zhang et al. 2025) 0.175 0.366 0.760 0.610 16.22 0.721 0.744 - 0.577 28.64
Table S3: Quantitative evaulation for fMRI-to-video reconstruction on CC2017 (Wen et al. 2018).
Method High-Level Low-Level
2-way 50-way BASIC-L MMR BASIC-H
MinD-Video (Chen, Qing, and Zhou 2023) 0.792 0.172 28.63 0.660 15.25
NeuroClips (Gong et al. 2024) 0.808 0.195 45.12 0.687 23.52
DecoFuse (Li et al. 2025) 0.824 0.208 29.03 0.706 15.31
Table S4: Quantitative evaluation for EEG-to-video reconstruction on SEED-DV (Liu et al. 2024b).
Method High-level Low-level
2-way 40-way BASIC-H 2-way 40-way SSIM BASIC-L
EEG2Video (Liu et al. 2024b) 0.798 0.159 49.54 0.774 0.138 0.256 20.54
Table S5: Quantitative evaluation for EEG-to-3D reconstruction on EEG-3D (Guo et al. 2025).
Method Average Top-1 of 5 Samples BASIC
2-way, Top-1 10-way, Top-3 2-way, Top-1 10-way, Top-3 BASIC-L BASIC-H
Neuro-3D (Guo et al. 2025) 55.81 35.89 72.08 57.64 9.69 23.08
Table S6: Quantitative evaluation for fMRI-to-3D reconstruction on fMRI-3D (Gao et al. 2024, 2025).
Methods Dataset Semantic-Level Structure-Level Textural-Level BASIC
[…]
MinD-3D (Gao et al. 2024) fMRI-Shape 0.828 0.459 3.157 1.742 3.833 0.306 32.81 0.674 14.72 30.95
MinD-3D++ (Gao et al. 2025) 0.887 0.616 3.025 1.635 3.672 0.234 34.09 0.763 12.62 35.08
MinD-3D (Gao et al. 2024) fMRI-Objaverse 0.793 0.427 4.304 2.142 5.323 0.544 31.09 0.724 N/A N/A
MinD-3D++ (Gao et al. 2025) 0.894 0.618 3.325 1.779 4.073 0.343 33.64 0.808 N/A N/A
N/A means reconstruction results on fMRI-Objaverse dataset are unavailable.
Appendix C Details on BASIC
This section complements the method section in the main paer with additional implementation details and qualitative results.
Structured Entity Extraction from Captions
Structured Descriptive Captioning.
We adopt LLaVA-1.6-13B (Liu et al. 2023) as our default MLLM, considering the trade-off between computational efficiency and captioning accuracy. Recent studies on detailed captioning evaluation (Lu et al. 2025; Dong et al. 2024) suggest that this configuration achieves top-tier performance in long image description, comparable to human judgments. For prompting, we adopt a simple instruction— “Describe the image in detail”—instead of crafting elaborate prompts that explicitly reflect evaluation dimensions outlined in Table 1. Preliminary experiments indicate that this simple prompt is sufficient to elicit detailed captions covering the desired evaluation dimensions when using LLaVA (Liu et al. 2023) . This may be attributed to LLaVA’s training corpus, which likely includes similarly phrased instructions for detailed image captioning. Further discussions on the qualitative and quantitative effects of different prompts and MLLMs are presented in the following sections.
Semantic Parsing and Matching.
This step extracts and matches three core visual elements, objects, attributes, and relationships, from candidate captions. The extracted entities are then evaluated using precision, recall, and F1 score. To facilitate intuitive understanding, we further decompose this process and demonstrate each component in detail. Consider the following two captions as the reference and candidate descriptions, respectively:
  • Reference
    ”A peaceful beach with soft white sand stretching along the coastline, where turquoise ocean waves gently roll onto the shore. Several people are sunbathing near the water while others are playing volleyball in the distance.”
  • Candidate
    ”A tropical island beach with lush palm trees swaying in the breeze, and the bright blue sea sparkling under the sun. Some people are lounging under umbrellas by the water, while a group of friends is playing volleyball near the edge of the beach.”
The objects, attributes, and relationships extracted from the captions are listed below. The attributes associated with the same object across different sentences, as well as spatial relationships - which are less intuitive and require further analysis to interpret the positional context - can also be extracted.
  • Reference
    objects: volleyball, sand, water, beach, people, shore
    attributes: beach: peaceful; sand: soft, white; people: sunbathing; wave: turquoise
    relations: (people, play, volleyball); (people, sunbath, water); (people, play near, water)
  • Candidate
    objects: water, people, island, beach, umbrella, sea
    attributes: island: tropical; sea: bright blue; palm tree: lush
    relations: (umbrella, by, water); (people, lounge under, umbrella); (people, on edge of, beach); (beach, on, island); (people, play, volleyball); (sea, under, sun)
Given that substances extracted from ground truth and candidate captions often differ in wording, we apply a stepwise matching strategy: (i) exact matching, (ii) synonym matching, and (iii) cosine similarity of word embeddings to resolve remaining unmatched elements via semantic alignment.
The evaluation scores for objects, attributes, and relations are computed as follows: (a) Precision, defined as the fraction of correctly predicted items among all items mentioned in the candidate description; (b) Recall, defined as the proportion of correctly predicted items among all items in the reference; (c) F1 score, the harmonic mean of precision and recall.
$\text{Precision}=\frac{N(\text{Matched})}{N(\text{Candidate})},\quad\text{Recall}=\frac{N(\text{Matched})}{N(\text{Reference})},\quad\text{F1}=\frac{2\cdot\text{Precision}\cdot\text{Recall}}{\text{Precision}+\text{Recall}},$ (1)
where $N(\text{Matched})$ denotes the number of items in the candidate that correctly match those in the reference, $N(\text{Candidate})$ is the total number of items in the candidate, and $N(\text{Reference})$ is the total number of items in the reference.
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
foreground background component foreground background component
Figure S1: The visualization of multigranular segmentation. Reference images and structured annotations can be found in Fig. S2 and Table S8, respectively.
Multigranular Image Segmentation
This section provides additional implementation details and qualitative results to complement the main paper.
Semantic Instance Categorization.
Multigranular segmentation involves decomposing visual scenes into hierarchical levels of detail, from whole-object recognition to fine-grained part segmentation. We first identify multigranular semantic instances that serve as textual instructions for subsequent segmentation using a multimodal large language model (Liu et al. 2023). The full categorization prompt used for multigranular object identification is provided in Table S7. As illustrated in Table S8, each image is annotated with structured information, including foreground objects, their constituent parts, and relevant contextual background elements. This hierarchical annotation schema allows for fine-to-coarse segmentation and supports comprehensive evaluation of vision models in terms of object-level recognition and detailed part-level understanding across diverse scene contexts.
Progressive Granular Segmentation.
Using the structured annotations described above as natural language prompts, we apply Grounded-SAM2 (Ren et al. 2023) to produce multigranular segmentation masks. Prompts such as “zebra”, “tail”, or “legs” lead the model to localize and segment corresponding entities and component within an image. This enables model evaluation on fine-grained visual grounding, compositional understanding, and spatial reasoning. The resulting multigranular segmentation visualizations are shown in Fig. S1, with corresponding reference images displayed in Fig. S2 and structured annotations provided in Table S8. Taking the top-right three images of the airplane as an example, we present the salient foreground (airplane), background (sky, runway, grass, and trees), and the components of the airplane (fuselage, wings, engines, and tail), respectively.
The metrics are then calculated between the reconstruction and ground truth over a hierarchical structure: salient (foreground categories), binary (foreground and background), semantic (all distinct categories), instance (individual instances of all identified object categories), and part (sub-object components, mostly salient foreground objects). This structured evaluation facilitates systematic and fine-grained assessment of model performance across multiple levels of visual granularity.
Please note that semantic and instance in the context of BASIC-L evaluation are technical terms from the task of image segmentation, referring to semantic segmentation (labeling each pixel with a class) and instance segmentation (distinguishing individual object instances), and should not be confused with their broader conceptual meanings.
Table S7: The categorization prompt for multigranular object identification.
You are a visual understanding assistant. Given an input image, please analyze and describe it with a structured categorization. Each response should include: 1. Foreground Objects: List the main objects in the foreground and their semantic categories. 2. Background Elements: Describe elements in the background and their semantic categories. 3. Part-level Categories: For each foreground object, identify the visible parts in a general way (e.g., ”car” $\rightarrow$ ”wheels”, ”doors”). Provide your answer in a structured JSON format like this:
  "foreground_objects": [
      "object": "dog",
      "semantic_category": "animal",
      "parts": ["head", "legs", "tail", "fur"]
  "background_elements": [
      "element": "tree",
      "semantic_category": "plant"
      "element": "sky",
      "semantic_category": "natural"
Table S8: Structured scene annotations for four image instances. ID 0 to 3 corresponds to references in Fig. S2. Each scene is decomposed into foreground (F) objects (e.g., boats, airplane, zebra), their associated part-level components, and background (B) elements (e.g., trees, sky, runway) that represent the environmental context. This is from a standardized JSON format.
ID Type Details
0 F boats: hull, deck, mast, sail; dock: pilings, decking, floating pontoons, fenders
B water, trees, sky, shore
1 F airplane: fuselage, wings, engines, tail
B sky, runway, grass, trees
2 F sink: basin, faucet, pipes; toilet: tank, bowl, seat, handle; cabinet: door, drawer, shelf
B tiles, window, mirror, shower curtain, plants
3 F zebra: head, neck, torso, tail, legs
B grass, trees, sky
Appendix D Qualitative Comparative Analysis
This section presents additional qualitative results and analyses, including brain visual decoding reconstructions, detailed captions for both reconstructions and stimuli, parsed caption annotations, and ablations on using different MLLMs and prompts. The results are primarily based on methods applied to NSD (Allen et al. 2022) for a comprehensive and in-depth examination.
Brain Decoding Reconstruction
Fig. S2 presents qualitative comparisons of brain visual decoding methods on the NSD dataset, including the original visual stimuli and reconstructions from SDRecon (Takagi and Nishimoto 2023b), BrainDiffuser (Ozcelik and VanRullen 2023), MindEye (Scotti et al. 2023), DREAM (Xia et al. 2024b), MindEye2 (Scotti et al. 2024), MindBridge (Wang et al. 2024a), UMBRAE (Xia et al. 2024a), NeuroPictor (Huo et al. 2024), NeuroVLA (Shen et al. 2024), UniBrain (Wang et al. 2024b), STTM (Liu et al. 2025), MindTuner (Gong et al. 2025), and BrainGuard (Tian et al. 2025). These results encompass the majority of current brain visual decoding methods on NSD, providing a representative and comprehensive overview that reflects the latest advancements and trends in the field.
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Reference SDRecon BrainDiffuser MindEye DREAM MindEye2 MindBridge
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
UMBRAE NeuroPictor NeuroVLA UniBrain STTM MindTuner BrainGuard
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Reference SDRecon BrainDiffuser MindEye DREAM MindEye2 MindBridge
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
UMBRAE NeuroPictor NeuroVLA UniBrain STTM MindTuner BrainGuard
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Reference SDRecon BrainDiffuser MindEye DREAM MindEye2 MindBridge
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
UMBRAE NeuroPictor NeuroVLA UniBrain STTM MindTuner BrainGuard
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Reference SDRecon BrainDiffuser MindEye DREAM MindEye2 MindBridge
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
UMBRAE NeuroPictor NeuroVLA UniBrain STTM MindTuner BrainGuard
Figure S2: Comparison of brain decoding reconstructions and corresponding references from NSD.
Detailed Captioning Comparison
Below is a comparison of detailed captioning results based on reconstructions generated by different brain decoding methods on NSD. The visual stimuli from the NSD test set, along with the corresponding decoded reconstructions for each image, are provided as reference in Fig. S2. Due to page limitations, we only present the detailed captions for the first image here (“a boat at a lake dock”).
We highlight the objects, attributes, and relationships extracted from the captions using the pipeline introduced in the main paper. These provide a structured and interpretable representation of the decoded contents. It should be noted that the annotations in the original captions shown below are intended for intuitive understanding and differ slightly from the parsed captions used for final evaluation, which are stored as structured data. The actual parsed results are below.
  • Method
    Detailed Captioning
  • Reference
    The image depicts a serene lakeside scene. In the natural foreground, there is a calm body of water with a few ripples, suggesting a gentle breeze. On the left side of the image, there is a wooden dock extend into the water, with a small boat moored at the end. The dock appears to be made of wooden planks and is support by wooden wooden pylons. In the middle of the image, there is a large boat dock at a floating dock. This boat is white and has a blue stripe run along its side. It has a covered area with a roof, and there are windows visible on side of the area, indicating it may be a cabin cruiser or a houseboat. On the right side of the image, there is a grassy hillside with trees, suggesting a wooded area. The sky is overcast, with no visible sun or shadows, which gives the scene a diffused light. There are no visible texts in the image or distinctive brands. The style of the image is a straightforward, color photograph. The focus is on the boats and the surrounding natural environment, with no people or other significant objects in the immediate foreground.
  • SDRecon
    The image captures a bustling scene at an terminal. The terminal is characterized by a curved ceiling, adorned with a pattern of diagonal lines on the ceiling that add a sense of depth and dimension to the space. A large sign, emblazoned with the words on the sign—”FLIGHTS” and ”GATE”—hang from the ceiling, guiding travelers to their respective destinations. The terminal is abuzz with activity. Numerous people walking can be seen moving around, each engrossed in their own journey. Their movements create a dynamic tableau, add life and add energy to the otherwise static environment. The terminal is also home to a variety of scattered vehicles, including cars and buses, which are in the terminal alongside the people. These vehicles, along with the people, contribute to the overall sense of movement and activity in the terminal. Despite the monochrome palette of the image, the scene is rich in detail and texture. The interplay of light and shadow, the intricate patterns on the ceiling, and the varied elements within the terminal all come together to create a vibrant, engaging image. The image is a testament to the constant motion and energy that characterizes life at an airport terminal.
  • BrainDiffuser
    The image is a photograph depicting a serene scene by a body of water, which appears to be a lake or a calm river. The water is reflecting the blue of the sky, suggesting it is a clear day. On the right side of the image, there is a dock or pier extend into the water, with several boats moor along its length. The boats vary in size and color, with one being particularly large and blue. On the dock, there are a few visible people, some of whom are standing and others who might be seated or moving about. They wear casual attire, and their presence adds a sense of activity to the otherwise tranquil setting. In the background, there are trees and a building with a sloped roof, which could be part of a residential or waterfront area in the background. The sky is partly cloudy, with patches of blue visible, indicating fair weather conditions.
  • MindEye
    The image depicts a serene water scene with a large blue and white boat float on calm water. The boat has a white superstructure with a blue hull and is equipped with multiple levels of seating and a covered area on the upper deck. There are several people on the boat, suggesting it is in use. The sky is clear with a few clouds, indicating fair weather. In the background, there are other boats moor at a dock, and the shoreline is lined with greenery. The style of the image is a realistic photograph with natural colors and lighting. There are no visible texts or distinctive marks that provide additional context or information about the location or the boat’s identity.
  • DREAM
    The image captures a serene scene at a marina during sunset. The sky, painted in orange and pink serve as a beautiful backdrop to the tranquil water. The marina is filled with numerous docked and scattered boats of various sizes, all bobbing gently on the surface of the water. The boats are scattered across the image, some closer to the foreground, while others are further away, creating a sense of depth. The calm water reflects the vibrant colors of the sky, adding to the overall tranquility of the scene. The image exudes a sense of peace and tranquility, as if inviting the viewer to take a moment and appreciate the beauty of nature.
  • MindEye2
    The image depicts a serene scene on a body of water, possibly a lake or a calm river. There are several boats visible, all of which appear to be small and possibly used for recreational purposes. The boats are painted in various colors, including shades of blue, white, and brown. In the foreground, there is a boat with a blue hull and a white cabin. This boat is closer to the viewer and is the most detailed in the image. Behind it, there are two other boats. The one in the middle has a brown hull and a white cabin, while the one on the far right has a white hull and a brown cabin on the right side of the hull. The calm, still water reflects the light in a way that suggests it might be a sunny day. The background is filled with lush greenery, indicating that the location is likely near a forest or a heavily wooded area. There are no visible texts or distinctive markings that provide additional information about the location or the boats. The style of the image is a standard, non-stylized photograph with a focus on the natural scenery and the boats. The perspective is from a distance, allowing for a view of the boats in the photograph and the surrounding environment. The image does not contain any people or animals, and the focus is on the stillness of the water and the tranquility of the setting.
  • MindBridge
    The image captures a serene scene of a river, where several boats are docked. The river, a deep shade, blue, is nestled amidst a backdrop of lush green trees. The boats, varying in size and color, are docked along the riverbank, their hulls glistening under the sunlight. In the foreground, a white boat stands out, its yellow roof adding a pop of color. Further down the river, a blue boat with a white roof can be seen, its vibrant hue contrasting with the surrounding greenery. The river itself is dotted with several buoys, their bright colors standing out against the blue of the water. These buoys, along with the docked boats, suggest a well-regulated waterway. In the distance, beyond the river, a city skyline can be seen. The city’s buildings, though far away, add a touch of urban life to the otherwise natural landscape. Overall, the image presents a harmonious blend of nature and urban life, with the river serving as a tranquil waterway amidst the city’s bustling activity.
  • UMBRAE
    The image captures a serene scene on a river, where two boats are gently floating. The boat on the left is a traditional wooden brown boat, painted in a warm brown color. It’s adorned with a white roof, adding a touch of rustic charm. On the right, there’s a modern blue and white boat, its sleek design contrasting with the wooden boat’s rustic appeal. The boats are not alone in the water; they are accompanied of a few people who are enjoying their time on the river. In the distance, you can see a small island with a tall lighthouse standing tall. The lighthouse, a beacon of safety for sailors, adds a sense of depth and scale to the scene. Above, the sky is a clear blue, dotted with fluffy white clouds. The sun is shining brightly, casting a soft glow on the water and illuminating the scene with a soft light. Overall, the image paints a picture of a peaceful day on the river, with boats gently floating and people enjoying their time on the water.
  • NeuroPictor
    The image depicts a serene waterfront scene with two boats docked at a pier. The boat on the left is a white motorboat with a blue stripe on side of its side. It has a small cabin and is equipped with a white outboard motor. The boat on the right is a large vessel with a blue hull and a white cabin. It has a white outboard motor as well. Both boats are secure with ropes. The pier itself is constructed of wood and is situated in front of a backdrop of lush green trees. The water is calm, reflecting the clear blue sky above. There are no visible markings on the boats or the pier that provide additional information. The style of the image is a realistic photograph with natural lighting and colors.
  • NeuroVLA
    The image captures a vibrant scene at a race track. The track, a smooth gray surface, is bordered by a grassy area on one side and a line of trees on the other. A group of people, dressed in a variety of colors, are scattered around the track, some standing, others walk around the track.
    On the track, two individuals are engaged in a race. One of them is leading, their body leaning forward in anticipation. The other racer is close behind, their body language suggesting a fierce competition. In the background, a blue building stand against the greenery. A red and white striped awning extends from it, providing a splash of color to the scene. The building’s location suggests it might be a facility related to the race track. The image is a dynamic snapshot of a moment, fill with color. It’s a testament to the excitement and energy of racing.
  • UniBrain
    The image depicts a coastal scene with a sandy beach leading up to a body of water that appears to be a river or a lake. On the beach, there are several boats of different sizes, including a small motorboat and a large sailboat. The boats are moored and appear to be unoccupied. A few people are visible, but they are too far away to discern any specific details. On the right side of the image, there is a road or path run parallel to the water’s edge. Along this path, there are various vehicles parked, including a large truck with a trailer attached and a smaller truck. The vehicles are park in a line, suggesting they might be used for transporting goods or equipment. The sky is overcast, and the lighting in the image is soft, indicating either early morning or late afternoon. The vegetation along the water’s edge includes a mix of grasses and small trees. There are no visible texts or distinctive brands in the image. The style of the image is a standard, non-stylized photograph with a focus on the natural and man-made elements present in the scene.
  • STTM
    The image captures a serene scene at a marina. Dominating the foreground is a large body of water, its surface rippling gently. A blue boat, adorned with a white stripe, is seen glid across the water, heading towards the right side of the image. On the left side of the image, a wooden pier extend into the water. It’s a hub of activity with several boats dock at it. Among them, a white boat stands out, its vibrant color contrast with the surrounding structures. In the background, a cluster of trees can be seen, their lush green foliage adding a touch of nature to the scene. Above it all, the sky stretches out in a clear blue expanse, devoid of any clouds. The image beautifully encapsulates a day at the marina, filled with the tranquility of the water and the bustle of boating activity.
  • MindTuner
    The image shows a large, white boat with a blue roof and a red stripe on the side, floating on a body of water. The boat has multiple levels and appears to be a passenger or tourist vessel, as suggested by the presence of windows and what looks like a covered area on the upper deck. The boat is equipped with a mast and sails, indicating that it may be capable of sailing. The surrounding environment includes a forested area with lush green trees, and the water reflects the overcast sky above. The style of the image is a realistic photograph with natural lighting and colors.
  • BrainGuard
    The image depicts a serene harbor scene. There are several boats of varying sizes and colors moored at the dock. The water is calm, reflecting the clear blue sky above. The dock is made of wood and extends into the water, providing a place for the boats to rest. In the distance, you can see a lush green hillside. The boats are securely tied to the dock, suggesting that they are not in motion. The overall atmosphere of the image is peaceful and idyllic.
Parsed Descriptions
The parsed captions are stored in a structured dictionary format, containing information about objects, attributes, and relationships. The identified objects are stored in a list, with attributes mapped to each object through key-value pairs. Relationships between objects are encoded in a directional scene graph, where edges capture the directed interactions or spatial relationships.
  • Method
    Parsed Description
  • Reference
    objects: boat, water, pylon, roof, ripple, window, stripe, tree, dock, sky, hillside
    attributes: water: calm; boat: large, white, small; dock: wooden, floating; pylon: wooden; stripe: blue; area: covered; hillside: grassy; sky: overcast, diffused; photograph: color, straightforward; foreground: natural
    relations: (window, on side of, area); (dock, extend into, water); (boat, dock at, dock); (hillside, with, tree); (boat, in, foreground); (water, with, ripple); (area, with, roof); (stripe, run along, boat); (dock, support by, pylon)
  • SDRecon
    objects: line, vehicle, word, terminal, people, ceiling, sign
    attributes: ceiling: curved; line: diagonal; sign: large; people: walking; tableau: dynamic; vehicle: scattered; pattern: intricate; terminal: vibrant, engaging
    relations: (word, on, sign); (tableau, add, energy); (vehicle, alongside, people); (tableau, add, life); (vehicle, in, terminal); (terminal, have, ceiling); (scene, have, detail); (sign, hang from, ceiling); (pattern, on, ceiling); (scene, have, texture); (terminal, have, vehicle)
  • BrainDiffuser
    objects: water, building, people, boat, dock, river, sky, roof
    attributes: river: calm; scene: serene; sky: blue, partly cloudy, clear; boat: blue, large; people: visible, standing; attire: casual; area: waterfront, part, residential; roof: sloped; photograph: naturalistic; element: human-made; environment: natural
    relations: (people, wear, attire); (boat, moor along, water); (building, in, background); (dock, extend into, water); (scene, by, water); (people, on, dock); (building, with, roof)
  • MindEye
    objects: sky, shoreline, water, boat, superstructure, people, cloud, dock, hull, seating, greenery
    attributes: boat: blue, large, white; water: calm; hull: blue; superstructure: white; sky: clear; photograph: realistic
    relations: (superstructure, with, hull); (boat, have, seating); (boat, moor at, dock); (people, on, boat); (sky, with, cloud); (greenery, line, shoreline); (boat, have, hull); (boat, float on, water)
  • DREAM
    objects: boat, water, marina, peace, sky
    attributes: scene: serene; sky: vibrant, pink, orange; water: tranquil, calm; boat: docked, scattered
    relations: (image, exude, tranquility); (boat, on, surface); (image, exude, peace); (sky, serve, water); (marina, fill with, boat); (sky, reflect in, water); (water, have, surface); (scene, at, marina)
  • MindEye2
    objects: boat, water, cabin, hull, light, greenery
    attributes: scene: serene; boat: detailed, blue, small, white, painted; hull: blue, white, brown; cabin: white, brown; water: still, calm, tranquil; greenery: lush; photograph: non-stylized, standard
    relations: (boat, in, distance); (scene, on, water); (water, reflect, light); (background, fill with, greenery); (boat, in, photograph); (boat, have, hull); (cabin, on the right side of, hull)
  • MindBridge
    objects: boat, riverbank, water, roof, city, tree, waterway, buoy, hull, building, river, greenery, skyline
    attributes: river: blue, tranquil, deep shade; tree: lush, green; hull: glistening; boat: blue, white, docked; roof: white, yellow; buoy: bright; water: blue
    relations: (boat, dock along, riverbank); (buoy, along, boat); (boat, with, roof); (greenery, surround, boat); (city, have, building); (boat, dock in, river); (city, have, activity); (river, have, buoy); (boat, have, roof); (river, in, city); (river, serve as, waterway); (boat, have, hull); (city, have, skyline)
  • UMBRAE
    objects: sky, boat, water, roof, island, cloud, lighthouse, sun, light, people, river
    attributes: boat: warm, brown, blue, wooden, modern, rustic, white; charm: rustic; roof: white; design: sleek; island: small; lighthouse: beacon, tall, safety; cloud: fluffy, white; sky: blue, clear; light: soft; sun: shining, bright; day: day, peaceful
    relations: (people, on, water); (boat, float on, river); (roof, add, charm); (glow, on, water); (people, on, river); (island, with, lighthouse); (sun, cast, glow); (design, on the right side of, boat); (boat, in, water); (boat, float on, water); (boat, accompanie of, people)
  • NeuroPictor
    objects: pier, boat, water, motorboat, cabin, lighting, tree, stripe, hull, motor, sky, rope
    attributes: motorboat: white; stripe: blue; cabin: small, white; motor: outboard, white; boat: large; hull: blue; pier: wooden; tree: lush, green; sky: blue, clear; water: calm; lighting: natural; photograph: realistic
    relations: (marking, on, boat); (boat, secure with, rope); (boat, dock at, pier); (cabin, have, motor); (stripe, on side of, motorboat); (photograph, have, lighting); (boat, have, cabin); (pier, in front of, tree); (boat, on, pier); (boat, have, hull)
  • NeuroVLA
    objects: building, people, line, greenery, racer, tree, body, race track
    attributes: scene: vibrant, dynamic; area: grassy; track: smooth; body: leaning; building: blue
    relations: (individual, on, track); (scene, fill with, color); (tree, in, line); (tree, on side of, track); (racer, have, body language); (people, walk around, track); (scene, at, race track); (area, border, track); (building, stand against, greenery); (building, link to, race track)
  • UniBrain
    objects: sailboat, water, path, road, people, trailer, boat, motorboat, line, beach, truck, vehicle, tree, sky, lighting, vegetation
    attributes: beach: sandy; motorboat: small; sailboat: large; boat: moored; people: visible, far away; truck: large, small; lighting: soft; sky: overcast; tree: small; element: natural, man-made; photograph: non-stylized, standard
    relations: (vehicle, have, trailer); (vegetation, have, tree); (road, on the right side of, image); (motorboat, on, beach); (vegetation, on edge of, water); (beach, lead up to, water); (vehicle, park in, line); (vehicle, park along, path); (text, in, image); (road, run parallel to, edge)
  • STTM
    objects: pier, boat, water, marina, hub, tree, stripe, foliage, sky
    attributes: scene: serene; surface: rippling; boat: vibrant, blue, white; stripe: white; pier: wooden; foliage: lush, green; sky: blue, clear; water: tranquil
    relations: (tree, have, foliage); (boat, glid across, water); (pier, extend into, water); (boat, dock at, hub); (boat, have, stripe); (marina, fill with, activity); (boat, contrast with, structure); (marina, fill with, water); (tree, in, background); (water, have, surface)
  • MindTuner
    objects: water, boat, mast, roof, window, tree, stripe, deck, sky, sail, lighting
    attributes: boat: large, white, passenger; roof: blue; stripe: red; area: green, covered, forested; deck: upper; tree: lush; sky: overcast; lighting: natural; photograph: realistic
    relations: (area, with, tree); (boat, have, sail); (boat, have, mast); (photograph, have, lighting); (window, on, deck); (stripe, on side of, boat); (text, on, boat); (boat, have, window); (boat, float on, water)
  • BrainGuard
    objects: water, boat, harbor, dock, wood, sky, hillside
    attributes: boat: colors, varying sizes; sky: blue, clear; water: calm; hillside: lush, green; atmosphere: idyllic, peaceful
    relations: (boat, at, dock); (dock, make of, wood); (boat, tie to, dock); (dock, extend into, water)
Captioning with Different MLLMs
The following two subsections present ablation examples evaluating the impact of different MLLMs and prompting strategies on detailed captioning. For clarity and brevity, we report only the parsed descriptions of objectsattributes, and relationships in the structured format. Despite variations in the generated captions across different MLLMs and prompts, the evaluation scores and method ranking remain consistent, as shown in Fig. 3, with more detailed scores in the quantitative evaluation (see  Table S9 and Table S10 ).
  • MLLM
    Parsed Description
  • LLaVA-1.5-7B
    objects: lake, people, boat, tree, car, person, water, pier
    attributes: boat: small, large, few; people: scattered; car: parked
    relations: (pier, on, lake); (car, park on the right side of, scene); (boat, in, center); (person, stand on edge of, water); (boat, in, background); (people, near, boat); (tree, surround, boat)
  • LLaVA-1.5-13B
    objects: dock, ripple, hillside, stripe, boat, tree, sky, water, roof, pylon, window
    attributes: water: calm; boat: large, white, small; dock: wooden, floating; pylon: wooden; stripe: blue; area: covered; hillside: grassy; sky: diffused, overcast; photograph: straightforward, color; foreground: natural
    relations: (stripe, run along, boat); (window, on side of, area); (hillside, with, tree); (dock, support by, pylon); (dock, extend into, water); (boat, in, foreground); (boat, dock at, dock); (area, with, roof); (water, with, ripple)
  • LLaVA-1.6-7B
    objects: awning, marina, boat, tree, island, water, pier
    attributes: scene: serene; boat: black, white, back; tree: green, lush; island: spotted, small; water: calm; view: picturesque
    relations: (boat, dock at, pier); (island, in, marina); (boat, have, awning); (boat, in, marina); (marina, nest among, tree)
  • LLaVA-1.6-13B
    objects: boat, water, pylon, roof, ripple, window, stripe, tree, dock, sky, hillside
    attributes: water: calm; boat: large, white, small; dock: wooden, floating; pylon: wooden; stripe: blue; area: covered; hillside: grassy; sky: overcast, diffused; photograph: color, straightforward; foreground: natural
    relations: (window, on side of, area); (dock, extend into, water); (boat, dock at, dock); (hillside, with, tree); (boat, in, foreground); (water, with, ripple); (area, with, roof); (stripe, run along, boat); (dock, support by, pylon)
  • LLaVA-1.6-34B
    objects: tree, boat, hillside, dock, pier, stripe, waterfront, sky, water
    attributes: waterfront: serene; pier: wooden; boat: small, blue, white; stripe: blue; area: covered; water: calm; hillside: lush, green, verdant; sky: gray, overcast; dock: peaceful; tree: green; object: relative; landscape: natural; scene: waterfront, realistic
    relations: (stripe, on side of, boat); (boat, have, stripe); (boat, in, image); (tree, cover, hillside); (object, dock at, pier); (boat, dock at, pier); (boat, on, water); (pier, extend into, water); (boat, against, sky); (boat, on, dock)
  • GPT-4o
    objects: yacht, lake, shoreline, forest, dock, boat, houseboat, hill, water, pier, shore
    attributes: forest: green, dense; lake: calm; boat: large, small; motor yacht: white; top: flat, sleek; dock: metal, wooden; water: windless, calm; hill: forested; scene: tranquil, scenic
    relations: (boat, dock at, pier); (forest, surround, lake); (hill, in, background); (dock, connect to, boat); (dock, anchor to, shore); (forest, line, shoreline); (yacht, have, top); (boat, on, lake)
Captioning with Different Prompts
This section presents a comparison of different description prompts and their effects on the detailed caption generation, using LLaVA-1.6-7B. Specifically, we compare three description prompts: (1) “Describe the image in detail.”, (2) “Describe the image in detail, focusing on visible objects.”, and (3) “Describe the image in detail, focusing on visible objects and their relationships.” Below are the parsed descriptions of objects, attributes, and relationships from the obtained captions.
  • Prompt
    Parsed Description
  • ID #1
    objects: awning, marina, boat, tree, island, water, pier
    attributes: scene: serene; boat: black, white, back; tree: green, lush; island: spotted, small; water: calm; view: picturesque
    relations: (boat, dock at, pier); (island, in, marina); (boat, have, awning); (boat, in, marina); (marina, nest among, tree)
  • ID #2
    objects: lake, greenery, marina, sky, boat, tree, island, water, hull, pier
    attributes: boat: large, white, small; design: sleek; pier: wooden; water: calm, blue; hull: white; tree: green, lush; island: small; sky: clear, blue; marvel: man-made
    relations: (tree, on either side of, lake); (boat, in, water); (boat, dock at, pier); (marvel, in, setting); (island, punctuate, horizon); (boat, have, hull); (island, with, greenery); (pier, extend into, water); (lake, nest among, tree); (tree, in, scene); (boat, have, design); (boat, mirror, boat)
  • ID #3
    objects: lake, light, marina, sky, boat, foliage, tree, water, pier
    attributes: boat: modern, large, white, blue; design: sleek; pier: weathered, wooden; water: calm; tranquility: broken; form: small; foliage: green; tree: lush; light: soft; scene: peaceful; sky: overcast; marina: beautiful
    relations: (light, over, scene); (form, scatter across, water); (pier, extend into, water); (lake, nest among, tree); (boat, have, design); (boat, in, background); (pier, contrast with, boat); (sky, have, light); (marina, have, backdrop); (lake, have, foliage)
Comparison of Multigranular Segmentation Results
Fig. S3 shows the comparison of multigranular segmentation results between reference images and reconstructions. The segmentations are visualized at multiple levels, including semantic, instance, and part granularity. Unlike previous metrics that evaluate only low-level consistency with the reference using a single black-box score, our segmentation-based multigranular metric provides more interpretable insights. IoU and AP scores are higher when the reconstructions are structurally aligned with the reference. Taking examples in Fig. S3, reconstructions where an airplane or zebra appears in the correct position and orientation receive higher scores—a property that holds across different granularities. For instance, BrainDiffuser (Ozcelik and VanRullen 2023) yields lower scores due to incorrect airplane orientation, while MindEye (Scotti et al. 2023) experiences score drops resulting from perspective shifts caused by camera zoom. Overall, most methods produce reconstructions that preserve the high-level semantic structure of the original images, as reflected in the alignment of major object categories. However, finer-grained details, such as part boundaries or object instances in crowded scenes, are often less accurately recovered, suggesting limitations in capturing precise spatial layouts and textures.
Reference Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
BrainDiffuser Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
MindEye Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
MindEye2 Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
NeuroPictor Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
DREAM Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
UMBRAE Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
BrainGuard Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
foreground background component foreground background component
Figure S3: Comparison of multigranular segmentations between reference images and reconstructions.
Table S9: Evaluation results using different prompt strategies. Three prompts are considered: (1) “Describe the image in detail.”, (2) “Describe the image in detail, focusing on visible objects.”, and (3) “Describe the image in detail, focusing on visible objects and their relationships.” The evaluation scores remain stable and the ranking of methods is consistent despite numerical fluctuations across prompts, indicating that our metric is robust to moderate prompt variations.
Method Object Attribute Relation BASIC-H Rank
P R F1 P R F1 P R F1
Prompt #1 BrainDiffuser 57.68 58.61 57.77 14.70 47.55 20.97 42.57 43.50 42.69 40.04 7
UMBRAE 61.12 62.48 61.37 18.17 53.93 25.61 47.47 48.96 47.84 44.36 6
NeuroPictor 62.31 62.21 61.68 18.38 53.63 25.43 48.62 49.49 48.67 44.58 5
MindEye 62.76 60.86 61.27 19.87 53.13 27.11 49.09 48.01 48.17 44.98 4
MindEye2 62.09 63.39 62.21 18.93 53.30 26.10 49.15 50.36 49.40 45.20 3
BrainGuard 63.29 63.18 62.68 19.69 54.61 27.18 50.34 51.75 50.62 46.07 2
DREAM 64.97 64.27 63.94 19.52 52.98 26.94 52.73 54.52 53.18 46.99 1
Prompt #2 BrainDiffuser 57.75 58.42 57.57 13.77 48.75 20.03 43.11 43.91 43.19 39.68 7
UMBRAE 61.81 62.94 61.89 17.81 52.95 25.07 48.67 48.90 48.46 44.48 6
NeuroPictor 62.58 61.63 61.43 18.67 51.34 25.74 49.50 48.98 48.88 44.65 5
MindEye 62.99 63.27 62.60 18.46 50.78 25.43 50.53 50.20 50.03 45.22 4
MindEye2 63.45 61.59 61.99 19.42 51.99 26.44 50.32 48.95 49.27 45.23 3
BrainGuard 63.67 62.72 62.65 19.80 54.07 27.26 51.08 50.91 50.62 46.09 2
DREAM 64.91 63.83 63.55 19.45 52.58 26.78 52.97 53.47 52.80 46.69 1
Prompt #3 BrainDiffuser 57.19 58.69 57.48 13.11 46.73 19.07 44.35 45.24 44.53 39.53 7
UMBRAE 61.06 62.21 61.17 17.42 52.04 24.40 49.69 49.78 49.48 44.12 6
NeuroPictor 62.57 61.46 61.34 17.40 50.27 24.13 50.86 50.39 50.32 44.25 5
MindEye 62.45 60.87 61.09 18.27 50.55 25.15 51.20 49.81 50.17 44.53 4
MindEye2 62.49 62.68 61.94 17.93 50.59 24.74 51.12 51.33 50.94 44.86 3
BrainGuard 63.75 62.40 62.46 19.21 52.17 26.43 52.66 52.14 52.12 45.98 2
DREAM 64.59 63.88 63.39 19.32 51.77 26.57 54.60 54.62 54.28 46.84 1
Appendix E Quantitative Comparative Analysis
Evaluating NSD with the BASIC
This section presents a performance comparison of visual decoding methods on NSD (Allen et al. 2022) using the proposed BASIC metrics, which capture both structural (BASIC-L) and semantic (BASIC-H) alignment. The final BASIC score is computed as the average of BASIC-L and BASIC-H, reflecting the equal importance of low-level structural and high-level semantic alignment in brain visual decoding evaluation. Fig. S4 presents the results, with the values of the top-5 best performing methods annotated for each dimension. For structural alignment NeuroPictor (Huo et al. 2024), STTM (Liu et al. 2025), MindEye2 (Scotti et al. 2024), BrainGuard (Tian et al. 2025), and DREAM (Xia et al. 2024b) achieve the highest BASIC-L scores, reflecting better spatial reconstruction. In contrast, semantic alignment (BASIC-H) is best captured by NeuroVLA (Shen et al. 2024), DREAM (Xia et al. 2024b), STTM (Liu et al. 2025), BrainGuard (Tian et al. 2025), and MindTuner (Gong et al. 2025), indicating superior semantic preservation. NeuroPictor (Huo et al. 2024) ranks highest in the composite BASIC score, followed by STTM (Liu et al. 2025), BrainGuard (Tian et al. 2025), MindEye2 (Scotti et al. 2024), and DREAM (Xia et al. 2024b), reflecting more balanced performance across both perceptual dimensions. While NeuroVLA (Shen et al. 2024) excels in semantic similarity, it underperforms in structural precision, suggesting a trade-off between high-level conceptual encoding and spatial accuracy. These results demonstrate the complementary nature of BASIC-L and BASIC-H and underscore the importance of multigranular evaluation for benchmarking brain visual decoding.
Figure S4: BASIC results for NSD evaluation.
Stability across Evaluation Configurations
This section evaluates the robustness and consistency of our metric across multiple factors, including MLLM choices and prompt strategies in BASIC-H, as well as box and text threshold configurations in BASIC-L. The metric scores remain relatively stable and largely insensitive to such setup variations.
Table S9 presents the quantitative results obtained using different captioning prompts. Three prompt variants, same as above are considered: (1) “Describe the image in detail.”, (2) “Describe the image in detail, focusing on visible objects.”, and (3) “Describe the image in detail, focusing on visible objects and their relationships.” Despite slight differences in phrasing, the evaluation scores remain consistent across prompts, indicating that our metric is robust to moderate prompt variations and does not overly rely on specific wording. The MLLM used here is LLaVA-1.6-7B (Liu et al. 2023). The metric remains consistent and discriminative even with very close scores (refer to the demonstrated scores of UMBRAE (Xia et al. 2024a), NeuroPictor (Huo et al. 2024), MindEye (Scotti et al. 2023), and MindEye2 (Scotti et al. 2024)).
Table S10 presents the quantitative evaluation results obtained using different MLLMs, including LLaVA-1.5-7B, LLaVA-1.5-13B, LLaVA-1.6-7B, LLaVA-1.6-13B, LLaVA-1.6-34B, and GPT-4o. Despite variations in model version and scale, the overall metric scores remain stable, and the relative method rankings are mostly preserved. This consistency across MLLMs demonstrates the robustness of our evaluation framework with respect to the choice of captioning model.
Table S11 presents a cross-evaluation, where reconstruction captions from each method are generated using one MLLM and assessed against captions produced by different MLLMs as references. Since no human-ranked scores are available for relevance analysis (i.e., consistency with human expert judgments using metrics such as Pearson correlation, coefficient of determination, or Kendall rank correlation), and obtaining reliable human-annotated relative rankings for each method is challenging, we instead adopt a cross-model evaluation strategy. Specifically, we assess performance by varying the MLLM models used as the producer (i.e., generating captions for the reconstructed images) and the evaluator (generating reference captions). This cross-model setup reveals consistency across MLLMs, suggesting that reliance on human-written “ground truth” captions is not strictly necessary for evaluation. Table S12 presents the detailed subscores for BrainDiffuser in the cross-model evaluation.
Table S13 summarizes the effect of varying box and text thresholds on the multigranular segmentation evaluation. While threshold changes influence individual sub-scores, the relative rankings of methods remain consistent within each configuration. The final BASIC-L score demonstrates notable stability across all threshold settings, despite fluctuations in component metrics.
Table S10: Evaluation results with different MLLM choices. The evaluation scores remain stable and the method ranking is mostly consistent despite numerical fluctuations across MLLM choices.
Method Object Attribute Relation BASIC-H Rank
P R F1 P R F1 P R F1
LLaVA-1.5-7B BrainDiffuser 56.96 59.04 57.26 11.84 23.65 14.60 45.32 46.76 45.65 37.87 7
UMBRAE 62.51 62.94 61.65 15.04 30.50 18.70 52.68 52.55 52.10 42.56 6
NeuroPictor 63.79 61.94 61.71 16.26 31.12 19.88 53.40 52.70 52.59 43.15 5
MindEye 63.93 61.98 61.88 16.08 32.43 20.16 53.51 52.35 52.41 43.30 4
MindEye2 64.17 63.62 62.71 17.51 34.49 21.58 53.90 53.81 53.39 44.39 3
BrainGuard 65.26 64.84 64.04 18.36 35.71 22.72 55.85 55.22 55.00 45.70 2
DREAM 68.76 65.62 65.54 18.52 36.81 23.06 60.02 59.18 59.01 47.24 1
LLaVA-1.5-13B BrainDiffuser 57.34 59.53 57.66 11.66 23.69 14.26 45.45 47.18 45.90 37.95 7
UMBRAE 62.32 63.20 61.72 15.42 31.11 19.00 52.34 52.64 52.05 42.70 6
NeuroPictor 63.97 62.91 62.32 16.17 32.08 19.99 52.98 53.20 52.66 43.46 4
MindEye 63.46 61.90 61.59 16.51 33.25 20.49 53.40 52.06 52.29 43.29 5
MindEye2 63.78 64.63 63.12 17.30 33.97 21.24 53.45 54.03 53.32 44.41 3
BrainGuard 65.21 64.75 63.85 18.52 35.15 22.63 55.59 55.90 55.30 45.65 2
DREAM 69.06 66.02 66.01 20.53 40.73 25.34 60.17 59.15 59.17 48.37 1
LLaVA-1.6-7B BrainDiffuser 57.68 58.61 57.77 14.70 47.55 20.97 42.57 43.50 42.69 40.04 7
UMBRAE 61.12 62.48 61.37 18.17 53.93 25.61 47.47 48.96 47.84 44.36 6
NeuroPictor 62.31 62.21 61.68 18.38 53.63 25.43 48.62 49.49 48.67 44.58 5
MindEye 62.76 60.86 61.27 19.87 53.13 27.11 49.09 48.01 48.17 44.98 4
MindEye2 62.09 63.39 62.21 18.93 53.30 26.10 49.15 50.36 49.40 45.20 3
BrainGuard 63.29 63.18 62.68 19.69 54.61 27.18 50.34 51.75 50.62 46.07 2
DREAM 64.97 64.27 63.94 19.52 52.98 26.94 52.73 54.52 53.18 46.99 1
LLaVA-1.6-13B BrainDiffuser 57.86 59.09 58.07 13.31 45.83 19.37 43.20 44.41 43.49 39.68 7
UMBRAE 61.59 62.09 61.32 17.50 50.64 24.39 49.08 49.08 48.75 44.03 6
NeuroPictor 62.99 61.06 61.38 17.88 50.00 24.60 49.62 49.08 48.98 44.19 5
MindEye 62.95 60.65 61.27 18.16 51.19 25.09 49.98 48.42 48.84 44.31 4
MindEye2 62.57 62.13 61.73 17.87 50.14 24.73 49.72 49.17 49.07 44.39 3
BrainGuard 63.64 62.37 62.44 18.67 52.41 25.86 51.25 50.72 50.60 45.44 2
DREAM 65.62 63.06 63.56 18.97 50.69 25.91 53.46 53.22 52.92 46.38 1
LLaVA-1.6-34B BrainDiffuser 57.55 58.70 57.67 13.43 46.59 19.50 42.98 44.47 43.40 39.55 7
UMBRAE 62.75 60.57 61.12 17.45 50.59 24.30 49.21 48.12 48.31 43.83 6
NeuroPictor 61.28 61.98 61.16 17.39 51.77 24.37 48.50 48.82 48.34 43.88 5
MindEye 62.85 60.96 61.23 17.60 50.84 24.33 48.91 48.66 48.45 43.91 4
MindEye2 62.35 62.24 61.75 17.28 49.96 24.15 49.20 49.13 48.82 44.13 3
BrainGuard 63.29 62.54 62.32 18.49 52.60 25.66 50.62 50.27 50.08 45.21 2
DREAM 65.30 63.13 63.48 18.82 50.78 25.67 52.48 52.53 52.12 46.09 1
GPT-4o BrainDiffuser 57.23 58.78 57.55 13.22 45.50 19.06 44.12 44.41 43.99 39.44 7
UMBRAE 62.04 60.49 60.68 16.72 48.84 23.26 50.39 47.78 48.75 43.32 6
NeuroPictor 60.76 61.79 60.74 16.94 49.13 23.60 49.64 48.67 48.87 43.51 4
MindEye 62.27 60.82 60.87 16.83 48.50 23.29 50.20 48.57 49.04 43.47 5
MindEye2 61.78 61.95 61.31 16.49 47.68 22.97 50.08 48.84 49.14 43.54 3
BrainGuard 62.90 62.24 61.96 18.34 51.29 25.22 51.76 49.96 50.54 44.98 2
DREAM 64.72 63.01 63.09 18.64 49.89 25.38 53.81 52.71 52.91 45.97 1
Table S11: Cross-model evaluation. Each method’s reconstruction captions (recon) are generated using one MLLM and evaluated against reference captions (stimuli) produced by another different MLLMs. This setup assesses the impact of using various MLLMs as the caption producer and the evaluator.
Stimuli Recon BrainDiffuser MindEye MindEye2 DREAM UMBRAE NeuroPictor BrainGuard
1.5-7B 1.5-7B 37.87 42.56 43.15 43.30 44.39 45.70 47.24
1.5-7B 1.5-13B 37.82 42.70 42.90 42.93 44.41 45.59 47.42
1.5-7B 1.6-7B 35.05 39.68 39.51 40.70 40.59 41.73 43.79
1.5-7B 1.6-13B 34.90 39.29 38.85 40.50 40.50 41.63 43.03
1.5-13B 1.5-7B 37.80 42.34 42.95 43.13 44.44 45.33 47.45
1.5-13B 1.5-13B 37.95 42.70 43.46 43.29 44.41 45.65 48.37
1.5-13B 1.6-7B 35.43 39.94 39.88 41.05 41.28 42.14 43.71
1.5-13B 1.6-13B 34.85 39.86 39.32 40.54 40.64 41.75 43.38
1.6-7B 1.5-7B 32.65 36.02 36.11 36.23 36.28 37.50 38.49
1.6-7B 1.5-13B 32.81 35.92 36.14 36.16 36.46 37.80 38.94
1.6-7B 1.6-7B 40.04 44.36 44.58 44.98 45.20 46.07 46.99
1.6-7B 1.6-13B 39.78 44.15 44.24 44.82 44.73 45.63 46.68
1.6-13B 1.5-7B 32.58 35.44 35.43 35.25 35.52 36.55 37.83
1.6-13B 1.5-13B 32.69 35.40 35.48 35.26 35.86 36.76 38.21
1.6-13B 1.6-7B 39.08 43.44 43.40 43.51 43.50 44.39 45.48
1.6-13B 1.6-13B 39.68 44.03 44.19 44.31 44.39 45.44 46.37
Table S12: Cross-model evaluation results for BrainDiffuser.
Setup Object Attribute Relation BASIC-H
Stimuli Reon P R F1 P R F1 P R F1
BrainDiffuser 1.5-7B 1.5-7B 56.96 59.04 57.26 11.84 23.65 14.60 45.32 46.75 45.65 37.87
1.5-7B 1.5-13B 57.14 58.77 57.20 11.86 23.88 14.65 45.08 46.54 45.40 37.82
1.5-7B 1.6-7B 57.81 49.81 52.96 11.97 21.48 14.21 42.82 39.86 40.93 35.05
1.5-7B 1.6-13B 58.40 48.89 52.63 11.95 22.23 14.33 43.42 38.75 40.59 34.90
1.5-13B 1.5-7B 56.82 59.46 57.37 11.53 23.66 14.21 45.42 47.10 45.84 37.80
1.5-13B 1.5-13B 57.34 59.53 57.66 11.66 23.69 14.26 45.45 47.18 45.90 37.95
1.5-13B 1.6-7B 57.97 50.36 53.41 12.25 22.27 14.56 42.93 40.27 41.21 35.43
1.5-13B 1.6-13B 58.65 49.43 53.13 11.55 20.93 13.55 43.59 39.19 40.90 34.85
1.6-7B 1.5-7B 48.88 59.95 53.24 5.15 25.94 8.17 38.26 43.67 40.41 32.65
1.6-7B 1.5-13B 49.23 60.02 53.41 5.30 26.95 8.40 38.33 43.63 40.43 32.81
1.6-7B 1.6-7B 57.68 58.61 57.77 14.70 47.55 20.97 42.57 43.50 42.69 40.04
1.6-7B 1.6-13B 58.29 57.58 57.59 14.56 45.41 20.65 43.07 42.38 42.41 39.78
1.6-13B 1.5-7B 48.25 60.82 53.10 4.79 28.37 7.82 38.10 45.36 41.05 32.58
1.6-13B 1.5-13B 48.49 60.99 53.29 4.91 27.94 7.94 38.07 45.19 40.96 32.69
1.6-13B 1.6-7B 56.44 59.24 57.39 12.91 47.22 18.89 41.84 44.54 42.83 39.08
1.6-13B 1.6-13B 57.86 59.09 58.07 13.31 45.83 19.37 43.19 44.41 43.49 39.68
Table S13: Evaluation results under different box and text thresholds.
Threshold Salient Binary Semantic Instance Part BASIC-L
Box Text IoU AP IoU AP IoU AP IoU AP IoU AP
BrainDiffuser 0.25 0.30 17.96 20.21 38.98 45.85 18.66 20.78 20.09 1.94 7.86 7.82 16.65
0.25 0.35 16.67 18.50 32.48 39.27 22.23 24.73 17.34 1.72 8.25 8.22 16.54
0.25 0.40 14.60 16.10 25.64 31.26 25.76 28.55 15.99 1.54 8.76 8.67 16.57
0.30 0.30 6.52 7.34 22.72 26.26 18.65 20.77 28.66 2.97 8.01 7.66 15.39
0.30 0.35 5.96 6.65 19.36 22.85 22.24 24.74 23.83 2.52 8.76 8.45 15.06
0.30 0.40 5.49 6.02 15.50 18.39 25.76 28.56 19.70 2.18 8.91 8.54 14.79
0.35 0.30 6.26 6.97 20.14 23.87 22.18 24.64 32.71 3.73 8.48 8.06 17.30
0.35 0.35 6.25 6.98 20.30 23.93 22.24 24.74 32.64 3.77 8.71 8.41 17.34
0.35 0.40 5.75 6.31 16.29 19.33 25.76 28.56 27.50 2.99 8.91 8.54 16.82
MindEye 0.25 0.30 19.20 22.79 45.53 55.17 18.73 21.94 20.36 1.99 7.49 8.26 17.03
0.25 0.35 18.27 21.49 39.75 49.61 21.64 25.08 18.01 1.90 7.47 8.14 16.89
0.25 0.40 16.65 19.60 32.86 41.91 25.27 29.03 16.76 1.73 7.83 8.41 17.07
0.30 0.30 7.55 8.76 26.09 30.97 18.72 21.94 28.12 3.04 7.74 8.14 15.53
0.30 0.35 7.15 8.34 22.87 27.74 21.64 25.08 24.38 2.80 7.63 7.93 15.17
0.30 0.40 6.43 7.40 19.00 23.39 25.27 29.03 20.63 2.45 8.58 8.65 15.12
0.35 0.30 7.23 8.35 22.96 28.04 21.51 24.96 31.56 3.97 7.84 8.05 17.00
0.35 0.35 7.36 8.58 23.61 28.59 21.65 25.09 31.75 4.12 7.78 8.08 17.11
0.35 0.40 6.62 7.62 19.62 24.11 25.28 29.04 27.90 3.42 8.57 8.65 16.99
UMBRAE 0.25 0.30 21.33 24.62 40.96 48.53 18.94 21.16 20.29 2.15 8.43 8.47 17.89
0.25 0.35 20.34 23.49 35.99 43.24 22.57 25.03 17.91 1.99 9.03 8.99 18.03
0.25 0.40 18.00 20.78 29.67 36.01 25.84 28.72 16.70 1.80 8.87 8.79 17.81
0.30 0.30 8.38 9.70 23.68 27.43 18.95 21.17 28.32 3.25 8.78 8.46 16.09
0.30 0.35 7.95 9.24 21.38 24.91 22.58 25.05 24.37 2.88 9.43 9.07 16.01
0.30 0.40 6.73 7.74 17.21 20.30 25.86 28.74 20.35 2.46 9.41 8.89 15.50
0.35 0.30 8.03 9.32 21.70 25.34 22.47 24.93 32.58 4.30 9.67 9.31 18.11
0.35 0.35 8.19 9.53 22.15 25.77 22.58 25.05 32.59 4.39 9.64 9.28 18.18
0.35 0.40 6.95 7.99 17.78 20.95 25.85 28.74 28.31 3.56 9.41 8.89 17.50
DREAM 0.25 0.30 23.62 26.10 46.03 57.11 21.15 24.13 21.41 2.32 9.22 8.79 19.57
0.25 0.35 22.74 24.72 40.18 50.40 24.92 27.69 18.88 2.23 10.43 9.88 19.86
0.25 0.40 19.69 21.10 32.22 41.52 27.37 29.81 17.34 2.03 10.46 9.74 19.17
0.30 0.30 8.83 9.95 24.72 30.18 21.14 24.12 29.79 3.52 9.80 9.10 17.34
0.30 0.35 8.54 9.36 21.72 26.76 24.91 27.68 25.78 3.22 11.03 10.10 17.44
0.30 0.40 7.60 8.15 17.70 21.94 27.39 29.83 21.51 2.80 10.83 9.77 16.67
0.35 0.30 8.61 9.47 21.66 26.96 24.62 27.42 32.95 4.52 10.89 9.89 19.15
0.35 0.35 8.72 9.56 22.23 27.36 24.91 27.68 33.22 4.70 11.01 10.08 19.35
0.35 0.40 7.76 8.32 17.68 22.40 27.39 29.83 28.99 3.74 10.83 9.78 18.59
BrainGuard 0.25 0.30 25.90 28.07 49.22 58.99 23.34 25.21 23.98 3.29 10.82 10.32 21.77
0.25 0.35 24.43 26.52 42.69 51.86 27.09 29.01 20.68 3.03 11.31 10.71 21.53
0.25 0.40 22.12 23.86 34.67 42.80 30.62 32.75 19.04 2.75 11.87 11.06 21.42
0.30 0.30 10.45 11.35 29.24 34.15 23.34 25.20 32.67 4.88 11.39 10.51 19.42
0.30 0.35 9.71 10.57 25.45 29.86 27.07 28.99 28.03 4.34 13.11 11.96 19.31
0.30 0.40 8.58 9.32 20.49 24.49 30.61 32.75 23.87 3.78 13.91 12.61 18.97
0.35 0.30 9.93 10.80 26.03 30.81 26.95 28.84 36.23 6.21 12.97 11.84 21.37
0.35 0.35 10.04 10.93 26.41 30.98 27.08 29.01 36.26 6.34 13.26 12.13 21.50
0.35 0.40 8.88 9.64 21.12 25.24 30.61 32.75 31.72 5.12 14.14 12.89 21.07
MindEye2 0.25 0.30 25.29 26.27 47.93 57.52 24.33 25.68 24.09 3.45 12.32 11.03 22.16
0.25 0.35 24.20 25.17 41.83 51.74 27.89 29.26 21.03 3.21 13.38 11.95 22.17
0.25 0.40 22.27 22.90 35.64 44.47 31.47 32.84 19.50 3.06 14.26 12.64 22.27
0.30 0.30 9.79 10.22 27.75 32.74 24.35 25.69 32.93 5.10 12.80 11.28 19.82
0.30 0.35 9.32 9.73 24.53 29.56 27.91 29.27 28.51 4.53 14.09 12.51 19.72
0.30 0.40 8.85 9.18 20.70 25.12 31.48 32.84 24.21 4.04 14.20 12.54 19.41
0.35 0.30 9.49 9.85 24.73 29.79 27.65 28.98 36.58 6.45 14.09 12.54 21.72
0.35 0.35 9.51 9.92 25.05 30.14 27.90 29.26 36.79 6.56 14.19 12.65 21.86
0.35 0.40 9.00 9.34 21.13 25.59 31.48 32.84 32.51 5.43 14.18 12.52 21.53
NeuroPictor 0.25 0.30 29.45 31.29 47.79 56.48 27.97 29.57 26.47 4.08 17.17 15.84 25.88
0.25 0.35 27.64 28.96 41.30 49.20 32.34 33.68 23.04 3.79 18.36 16.67 25.81
0.25 0.40 24.45 25.65 33.91 41.12 37.05 38.35 21.30 3.59 21.01 19.13 26.13
0.30 0.30 11.65 12.26 27.75 31.91 27.97 29.57 36.13 6.03 18.06 16.17 23.13
0.30 0.35 10.84 11.42 24.56 28.30 32.34 33.69 31.31 5.31 19.79 17.43 23.12
0.30 0.40 9.47 9.97 20.04 23.37 37.06 38.37 26.68 4.78 21.61 19.30 23.10
0.35 0.30 11.22 11.78 25.29 29.16 32.37 33.76 40.78 7.59 19.85 17.52 25.62
0.35 0.35 11.17 11.77 25.44 29.24 32.34 33.68 40.75 7.64 19.90 17.54 25.61
0.35 0.40 9.77 10.28 20.81 24.20 37.06 38.36 35.88 6.29 21.72 19.40 25.51
Appendix F Limitation and Broader Impact
Limitations.
Despite our efforts in developing a versatile and robust evaluation pipeline, there are still challenges in capturing nuances in brain visual decoding, such as exploring additional cognitive dimensions like emotional or attentional states, to enable a more holistic evaluation. This limitation stems from current datasets lacking sufficient diversity in cognitive states or subject demographics. Furthermore, our method targets recent decoding techniques that can generate clear and semantically relevant images. The aspect of image quality is therefore not considered a key criterion. This makes the framework less suitable for assessing early-stage visual coding work that often contains artifacts, blurriness, or difficult-to-recognize semantics. The reliance on large-scale MLLMs also presents challenges in terms of computational cost and susceptibility to hallucination.
Broader Impacts.
This presented framework provides a more detailed, interpretable, and standardized method for assessing visual decoding models. These enhanced evaluation capabilities facilitates a deeper understanding of how models interpret and reconstruct visual information, which is crucial for advancing brain-inspired models, brain-computer interfaces, and assistive technologies. To address privacy and security concerns, particularly those surrounding neurodata and potential misuse, it is essential to rigorously assess ethical implications and implement safeguards for participant data before applying the evaluation.